BIBLIOGRAPHY OF VASSILIEV INVARIANTS

DROR BAR-NATAN AND SERGEI DUZHIN

This bibliography is available electronically at http://www.pdmi.ras. ru/∼duzhin. The html version has some hyperlinks!

This bibliography was started in 1995 by Dror Bar-Natan who maintained it until 2005 when the project was taken over by Sergei Duzhin. Please send (coordinates of) papers, comments, hyperlinks etc. to my lastname at pdmi.ras.ru.

Contents

1. List of Additions 2 2. Electronic Addresses 5 3. Acknowledgement 12 4. References 12 4.1. References beginning with A 12 4.2. References beginning with B 13 4.3. References beginning with C 15 4.4. References beginning with D 17 4.5. References beginning with E 17 4.6. References beginning with F 18 4.7. References beginning with G 18 4.8. References beginning with H 21 4.9. References beginning with I 22 4.10. References beginning with J 22 4.11. References beginning with K 22 4.12. References beginning with L 24 4.13. References beginning with M 27 4.14. References beginning with N 29 4.15. References beginning with O 30 4.16. References beginning with P 30 4.17. References beginning with Q 31 4.18. References beginning with R 32 4.19. References beginning with S 32 4.20. References beginning with T 34 4.21. References beginning with U 35 4.22. References beginning with V 35 4.23. References beginning with W 37 4.24. References beginning with X 37 4.25. References beginning with Y 37 4.26. References beginning with Z 38

Date: Jul. 4, 2013; Entries: 0 .

1. List of Additions

Here’s a list of additions to this bibliography, beyond the first 100 papers. Note that if a paper stands under, say, July 2005, this does not mean that it was published around that time – this means that it was added to the bibliography in July 2005!

• July 2013 [V23]

• April 2011 [G49] [G50] [G51]

• January 2011 [G47] [G48]

• January 2010 [D17] [T2] [Z14]

• March 2009 [T18] [L9]

• October 2007 [H13] [B2] [S16] [A24] [M47] [B26] [K12]

• November 2006 [A6] [A7] [D19]

• October 2006 [Y2] [P11] [P10] [Z2]

• March 2006 [R4], [M24]

• January 2006 [C36], [V36], [P26], [B43].

• November 2005 [C40], [H10], [H12], [M23], [M44], [M49], [M50], [T14], [T17], [Y7].

• July 2005: [A2], [D16], [D18], [F7], [F12], [K32], [K53], [L12], [L18], [M3], [M15], [M41], [M42], [S52], [S34], [S53], [S37], [S19], [T1], [V11], [V12], [V13], [V14] , [V15], [V16], [V17], [V18], [V19], [V20], [V21], [V22].

• May 2005: [C1].

• March 2005: [T16], [H17], [G17], [C35].

• February 2005: [V35].

• January 2005: [C20], [M34].

• December 2004: [B1], [Y5], [M43].

• November 2004: [G32], [L30], [L31], [T15].

• October 2004: [M6], [K27].

• September 2004: [K65], [C30].

• August 2004: [Z1], [B15], [B16], [K9], [K4], [K5], [C25].

• July 2004: [E10], [P25], [M33], [P20].

• June 2004: [O11], [M20], [M5], [C34], [P5].

• May 2004: [S31], [T22], [T24], [K10], [K11].

• February 2004: [C37], [C33], [M21], [M22], [S12].

• January 2004: [A23], [V34].

• December 2003: [M25], [M32].

• November 2003: [T9].

• October 2003: [S10], [N15], [N16] [N4], [M4], [G12], [T5], [M12], [M13], [M14], [O10], [S11], [B25], [N2], [M9].

• September 2003: [M8], [G45].

• August 2003: [M2], [C6].

• May 2003: [V33], [K49], [B42].

• April 2003: [K43], [N9], [N10], [S1].

• February 2003: [F14].

• January 2002: [R6], [F8].

• November 2002: [M48].

• October 2002: [K60] [L29], [M10].

• September 2002: [C14], [H15], [M18], [C39], [B36], [J3], [K48].

• July 2002: [N14], [M7].

• June 2002: [G31] [G11].

• May 2002: [H4], [G33], [B29], [B30].

• April 2002: [L42], [M17], [B23], [V32], [Z4].

• March 2002: [G26], [S4].

• February 2002: [P23], [P4], [O9], [S18].

• January 2002: [G9], [G10], [S22], [O15], [R14], [F6], [P15].

• December 2001: [R2], [R3], [S9], [E1], [E2], [E3], [E4], [E5], [E6].

• November 2001: [M11], [N8], [K55], [T23], [B14].

• October 2001: [Q1], [C32], [G16], [P2], [P3], [G43] (English).

• June 2001: [L58], [T13], [R13], [T4].

• May 2001: [G15], [W9].

• April 2001: [Y4], [H23], [J7], [J2], [Y1], [O18], [J5], [J4], [O14], [N1], [L41].

• March 2001: [M26], [N13].

• February 2001: [G24], [G25].

• December 2000: [C13], [F5], [L54], [C38].

• November 2000: [L33], [D5].

• October 2000: [K44], [T12], [L53], [L49], [M35].

• September 2000: [V31], [S8], [V10].

• August 2000: [M52].

• July 2000: [B22], [L4].

• June 2000: [G23], [G34], [P27], [T8], [O1].

• May 2000: [H22], [C12], [V26], [K59], [H18], [A22].

• April 2000: [O2], [M31], [L28], [G13].

• March 2000: [T3], [M62], [B38], [G7], [G30], [G8], [F4].

• February 2000: [T6], [E8], [E9], [L3], [S6], [S7].

• January 2000: [V9], [L40], [K58].

• December 1999: [C3], [B3], [N11], [K64].

• November 1999: [K45], [C5], [P28], [O6], [H8], [S5], [D12], [D13], [P14], [V29], [V30], [L27], [R12], [O7], [O8], [F17].

• September 1999: [L39], [B33], [S49], [F9], [S50], [S51], [M53], [D11].

• July 1999: [O24], [R5], [S30], [M51], [G43], [C11], [C31].

• June 1999: [C10], [A19] [K40], [K41], [K42], [L17].

• May 1999: [G22], [A21], [M30], [L2], [M58], [T19]. [T20], [T21].

• April 1999: [E7].

• March 1999: [M1], [W8], [S32], [M57], [N12].

• February 1999: [M56], [P13].

• January 1999: [K57], [A18], [B28], [C9].

• December 1998: [M59], [G41], [G42], [P18], [P19], [H9], [L8], [J1], [S13], [S14], [P12], [M29].

• November 1998: [P1], [L38], [G6], [K25], [K26], [L32], [F3], [H14], [H21], [Z3], [K3], [C8].

• October 1998: [G44], [S47], [S48], [B4] [L57].

• September 1998: [K14], [K15], [C29].

• August 1998: [V6], [B21], [D10], [R1], [O16], [O17], [C7], [H11].

• July 1998: [M27], [S46], [L7], [S29], [M28].

• June 1998: [K37], [G2], [L37], [B27].

• May 1998: [H7], [H3], [S28], [F13],

• April 1998: [B35]. [K8], [P17], [K39], [K20].

• March 1998: [S43], [S44], [S45], [O19], [O20].

• February 1998: [L11], [W7], [L16], [A4], [A5], [T11], [M38], [M39], [M40].

• January 1998: [B20], [C28].

• December 1997: [L6], [C21], [W6].

• November 1997: [G1], [A3].

• October 1997: [Y6], [J6], [L5].

• September 1997: [C19], [V24], [W5], [Y3], [B40] [D4].

• August 1997: [G5], [H5].

• July 1997: [K7], [K2], [P6], [G14].

• June 1997: [B19], [L36], [M61], [S54], [K38], [S38], [S39], [S40], [S41], [S42].

• May 1997: [S15].

• April 1997: [K18], [G4], [G27], [K52], [L15], [A15].

• March 1997: [H2], [B18], [M19], [M65], [K31].

• February 1997: [V8], [M64], [K17].

• January 1997: [R15], [H6], [O5].

• December 1996: [K54], [B41], [B37], [O4], [M60], [G37], [D2], [D3], [F11].

• November 1996: [C18], [C16], [A17].

• October 1996: [S36], [H1], [L35].

• September 1996: [K51], [A16], [M16], [G21].

• August 1996: [K61], [P22], [P16], [K36].

• July 1996: [A14], [D1], [L19].

• June 1996: [T7].

• May 1996: [V2], [D8], [V3], [C15], [B46].

• April 1996: [A13], [B44], [L52].

• March 1996: [G20], [A10], [K56], [B32], [L34].

• February 1996: [K34], [F10], [S21], [K47].

• January 1996: [B24], [R10], [R11], [K13], [L51], [S35].

• December 1995: [O13], [W4], [S27], [M54], [M55], [K16], [L14].

• November 1995: [L1], [G29], [H20], [L26], [K62].

• October 1995: [L50], [F16].

• September 1995: [L56], [F15], [O22], [O23], [L10], [C26], [K6], [F2], [G19], [K35], [N5].

• August 1995: [K30], [G35], [G36], [L20], [G28], [B13].

• July 1995: [S17], [G40], [V25], [F1], [V28].

2. Electronic Addresses

Here are some authors’ email, http, and ftp addresses (notice that electronic addresses are constantly changing, so the list below cannot be reliable. Please let me know if you have any additions/corrections). As a precaution against spam, I publish email addresses in a disguised form (replacing dots by vlines and ats by stars) – please tell me if you want your address to be removed altogether.

A: (see page 12)

• Francesca Aicardi:

  aicardi*sissa|it

• Iain Aitchison:

• Peter M. Akhmetiev:

• Sergei Allyonov: allenov*list|ru

• Marcos Alvarez:

• Daniel Altschuler:

• Jørgen Ellegaard Andersen: andersen*imf|au|dk, http://home.imf.au.dk/andersen/, http://www.ctqm.au.dk.

• Eli Appleboim: eliap*ee|technion|ac|il

• V. I. Arnold: http://www.pdmi.ras.ru/˜ arnsem/Arnold/, http://www.institut.math.jussieu.fr/seminaires/singularites/pagearnold.html

• Nikos A. Askitas:

• Emmanuel Auclair:

http://www-fourier.ujf-grenoble.fr/Personnel/auclaire.html B: (see page 13)

• Sebastian Baader: baader*math-lab|unibas|ch

• Eric Babson:

• Roland Bacher:

• John Baez: http://math.ucr.edu/home/baez/README.html

• Dror Bar-Natan: drorbn*math|toronto|edu,

http://www.math.toronto.edu/˜ drorbn

• Anna Beliakova: http://www.math.unizh.ch/index.php?professur&key1=578

• Paolo Bellingeri: bellingeri*mail|dm|unipi|it

• Anna-Barbara Berger:

• Mitchell A. Berger: m.berger*ucl|ac|uk, http://www.ucl.ac.uk/ ucahmab/home.html

• Joan S. Birman: jb*math|columbia|edu,

http://www.math.columbia.edu/˜ jb/

• Anders Björner: http://www.math.kth.se/˜ bjorner/

• Béla Bollobás: http://www.msci.memphis.edu/faculty/bollobasb.html

• Raoul Bott [1923-2005]:

• David J. Broadhurst: D.Broadhurst*open|ac|uk,

http://physics.open.ac.uk/˜ dbroadhu

• Ryan Budney, http://guests.mpim-bonn.mpg.de/rybu/

• Doug Bullock:

• Urs Burri: C: (see page 15)

• Rutwig Campoamor-Stursberg

• Pierre Cartier

• Alberto S. Cattaneo: alberto|cattaneo*math|unizh|ch,

http://www.math.unizh.ch/asc

• Nafaa Chbili: http://www.math.titech.ac.jp/ chbili/

• Vladimir Chernov (=Tchernov): Vladimir|Chernov*dartmouth|edu,

http://www.math.dartmouth.edu/˜ chernov/

• Sergei Chmutov: chmutov*math|ohio-state|edu,

http://www.math.ohio-state.edu/˜ chmutov/.

• Tim D. Cochran: http://math.rice.edu/˜ cochran/

• Fred R. Cohen, http://www.math.rochester.edu/people/faculty/cohf/

• James Conant: jconant*math|utk|edu,

http://www.math.utk.edu/˜ jconant/home.html

• Paolo Cotta-Ramusino: http://wwwteor.mi.infn.it/users/cotta/

• Peter R. Cromwell D: (see page 17)

• John Dean:

• Corrado De Concini:

• Oliver T. Dasbach: http://www.math.lsu.edu/˜ kasten/

• Tetsuo Deguchi:

• Cayetano Di Bartolo

• Michel Domergue:

• Paul Donato:

• Sergei Duzhin: duzhin*pdmi|ras|ru,

http://www.pdmi.ras.ru/˜ duzhin/.

• Ivan Dynnikov: http://www.math.msu.su/˜ dynnikov/ E: (see page 17)

• Michael Eisermann: http://www-fourier.ujf-grenoble.fr/ eiserm/

• Tobias Ekholm: http://www.math.uu.se/˜ tobias/

• Benjamin Enriquez: enriquez*math|u-strasbg|fr

F: (see page 18) • Boris Fain:

• Roger Fenn: http://www.maths.sussex.ac.uk/Staff/RAF/

• Thomas Fiedler: fiedler*picard|ups-tlse|fr,

http://picard.ups-tlse.fr/homepage/fiedler.html

• José M. Figueroa-O’Farrill:

http://www.maths.ed.ac.uk/ jmf/

• Jonathan Fine:

• V.V.Fock: fock*math|brown|edu

• Laurent Freidel:

• Jürg Fröhlich:

• Charles Frohman:

• Dmitry B. Fuchs:

• Louis Funar: http://www-fourier.ujf-grenoble.fr/˜ funar/ G: (see page 18)

• Rodolfo Gambini

• Stavros Garoufalidis: stavros*math|gatech|edu,

http://www.math.gatech.edu/˜ stavros/

• Fabio Gavarini: gavarini*mat|uniroma2|it

• Nathan Geer: geer*math|gatech|edu http://www.math.gatech.edu/ geer/ • Sylvain Gervais: http://www.math.sciences.univ-nantes.fr/˜ gervais/

• Juan González-Meneses: http://deimos.us.es/˜ meneses/

• Victor Goryunov: goryunov*liv|ac|uk,

http://www.liv.ac.uk/PureMaths/members/V Goryunov.html, ftp://ftp.liv.ac.uk/pub/goryunov/

• Mikhail Goussarov [1958–1999]: http://euclid.pdmi.ras.ru/˜ goussar/, http://www.math.toronto.edu/˜ drorbn/Goussarov/

• Matías Graña: matiasg*dm|uba|ar,

http://mate.dm.uba.ar/˜ matiasg/

• Matt Greenwood:

• Jorge Griego H: (see page 21)

• Nathan Habegger: http://www.math.sciences.univ-nantes.fr/˜ habegger/,

  ftp://poincare.math.sciences.univ-nantes.fr/pub/preprint/
  alg-topo/Habegger/

• Kazuo Habiro: http://www.kurims.kyoto-u.ac.jp/˜ habiro/

• Ami Haviv:

• Nobuharu Hayashi:

• Laure Helme-Guizon:

• Vladimir Hinich: hinich*math|haifa|ac|il,

http://math.haifa.ac.il/hinich/WEB/hinich.htm

• Allen C. Hirshfeld:

• Hugh Nelson Howards: http://www.mthcsc.wfu.edu/˜ howards/

• Youngsik Huh:

• Michael Hutchings: http://math.berkeley.edu/˜ hutching/ I: (see page 22) J: (see page 22)

• Tadeusz Januszkiewicz: http://www.math.uni.wroc.pl/˜ tjan/

• Myeong-Ju Jeong:

• Gyo Taek Jin: http://knot.kaist.ac.kr/˜ trefoil/

• Jeff Johannes: http://www.geneseo.edu/ johannes/

• Peter M. Johnson: K: (see page 22)

• Efstratia Kalfagianni: http://www.math.msu.edu/˜ kalfagia/

• Taizo Kanenobu:

• Joanna Kania-Bartoszynska: http://diamond.idbsu.edu/˜ kania/

• Mikhail Kapranov: http://www.math.toronto.edu/˜ kapranov/

• Christian Kassel: http://www-irma.u-strasbg.fr/˜ kassel/

• Louis H. Kauffman: http://www.math.uic.edu/˜ kauffman/

• Akio Kawauchi: kawauchi*sci|osaka-cu|ac|jp,

http://www.sci.osaka-cu.ac.jp/ kawauchi/

• Mikhail Khovanov: http://www.math.columbia.edu/ khovanov

• Takashi Kimura:

• Paul Kirk: http://php.indiana.edu/˜ pkirk/

• Thomas Klöker

• Jan A. Kneissler: jan*kneissler|info,

http://www.math.uni-bonn.de/people/jk

• Ilya Kofman: ikofman*math|columbia|edu,

http://www.math.columbia.edu/˜ ikofman/

• Toshitake Kohno: http://mint.ms.u-tokyo.ac.jp/

• Maxim Kontsevich:

• Kirill Krasnov:

• Dirk Kreimer: http://math.bu.edu/people/dkreimer/

• Andrew Kricker: http://www.ntu.edu.sg/home/ajkricker/

• Greg Kuperberg: , http://www.math.ucdavis.edu/˜ greg

• Vitaliy Kurlin: vak26*yandex|ru,

http://www.geocities.com/vak26 L: (see page 24)

• J. M. F. Labastida: http://www-fp.usc.es/theory/labastida/

• Sergei K. Lando: lando*mccme|ru, http://www.mccme.ru/˜ lando/

• Jean Lannes

• Ruth Lawrence: ruthel*ma|huji|ac|il,

http://www.ma.huji.ac.il/˜ ruthel/

• Thang Q. T. Le: http://www.math.gatech.edu/˜ letu/index.html

• Jung Hoon Lee:

• Christine Lescop: lescop*ujf-grenoble|fr,

http://www-fourier.ujf-grenoble.fr/˜ lescop/

• Jerome Levine [1937–2006]: http://people.brandeis.edu/˜ levine

• Banghe Li:

• Jens Lieberum: http://swiss2.whosting.ch/jenslieb/

• Xiao-Song Lin [1957–2007]:: http://math.ucr.edu/˜ xl/

• Svante Linusson: http://www.math.kth.se/˜ linusson/

• Charles Livingston: http://php.indiana.edu/˜ livingst/home.html

• Martin Loebl: loebl*kam|ms|mff|cuni|cz,

http://www.ms.mff.cuni.cz/˜ loebl

• Riccardo Longoni: http://www.mat.uniroma1.it/ longoni

• John E. Luecke: http://rene.ma.utexas.edu:80/text/webpages/luecke.html M: (see page 27)

• Yuri B. Magarshak:

• Seth A. Major:

• Vassiliy O. Manturov:

• Julien Marché: http://www.institut.math.jussieu.fr/˜ marche/ • Marcos Mariño:

• Maruizio Martellini:

• Joao Faria Martins:

• Gregor Masbaum: http://www.math.jussieu.fr/˜ masbaum

• Gwénaël Massuyeau: http://www-irma.u-strasbg.fr/ massuyea/

• Josef Mattes: http://tyche.mat.univie.ac.at/˜ mattes/

• Sergei Matveev: matveev*csu|ru

• Darryl McCullough: http://www.math.ou.edu/˜ dmccullough/

• Michael McDaniel:

• Jean-Baptiste Meilhan: http://www.kurims.kyoto-u.ac.jp/ meilhan/ • Sergey A. Melikhov: melikhov*mi|ras|ru,

http://www.math.ufl.edu/ melikhov/

• Blake Mellor: http://myweb.lmu.edu/bmellor/

• Paul Melvin:

• Guowu Meng:

• Alexander B. Merkov: merx*mccme|ru

• Haruko Aida Miyazawa:

• Yasuyuki Miyazawa:

• Iain Moffatt: iainm*maths|warwick|ac|uk,

http://www.maths.warwick.ac.uk/˜ iainm/

• Hugh R. Morton: morton*liv|ac|uk,

http://www.liv.ac.uk/˜ su14/

• Daniel Moskovich: http://www.sumamathematica.com/

• Jacob Mostovoy: http://www.matcuer.unam.mx/˜ jacob/

• Hitoshi Murakami: starshea*tky3|3web|ne|jp,

http://www3.tky.3web.ne.jp/ starshea/index.htm

• Jun Murakami: http://www.f.waseda.jp/murakami/

• Kunio Murasugi N: (see page 29)

• Hiroaki Nakamura:

• Yasutaka Nakanishi: http://www.math.s.kobe-u.ac.jp/HOME/nakanisi/ • Gad Naot:

• N.A.Nekrasov:

• Ka Yi Ng:

• Hideaki Nishihara: http://homepage3.nifty.com/Nis/

• Steve Noble: mastsdn*brunel|ac|uk,

http://www.brunel.ac.uk/˜ mastsdn/,

http://people.brunel.ac.uk/ mastsdn

• Tahl Nowik: tahl*math|biu|ac|il,

http://www.math.biu.ac.il/˜ tahl/ O: (see page 30)

• Tomoshiro Ochiai:

• Eiji Ogasa:

• Tomotada Ohtsuki: tomotada*kurims|kyoto-u|ac|jp,

http://www.kurims.kyoto-u.ac.jp/˜ tomotada/

• Yoshiyuki Ohyama: http://lab.twcu.ac.jp/ohyama/ • Miyuki Okamoto:

• Ping-Zen Ong:

• Kent E. Orr: http://php.indiana.edu/ korr/

• Tetsuya Ozawa P: (see page 30)

• Paolo Papi: http://mercurio.mat.uniroma1.it/ricercatori/papi/ • Luis Paris: http://math.u-bourgogne.fr/topolog/paris/index.html

• Bertrand Patureau-Mirand: bertrand|patureau*univ-ubs|fr,

http://www.univ-ubs.fr/lmam/patureau/

• Chan-Young Park:

• Esther Pérez

• Roger Picken: rpicken*math|ist|utl|pt

• Sergey Piunikhin:

• Sylvain Poirier: http://spoirier.lautre.net/

• Michael Polyak: http://www.math.technion.ac.il/people/polyak/ • Viktor Prasolov:

http://www.mccme.ru/prasolov

• Claudio Procesi:

http://mercurio.mat.uniroma1.it/ordinari/procesi/home.html

• Jorge Pullin: pullin*lsu|edu

• Jozef H. Przytycki: przytyck*gwu|edu,

http://home.gwu.edu/˜ przytyck/ Q: (see page 31)

• Frank Quinn: http://www.math.vt.edu/people/quinn/ R: (see page 32)

• Richard Randell:

• Dusan Repovs: dusan.repovs*guest|arnes|si,

http://pef.pef.uni-lj.si/˜ dusanr/index.htm

• Nicolai Reshetikhin: http://math.berkeley.edu/ reshetik/

• Oliver M. Riordan:

• Justin Roberts: http://www.math.ucsd.edu/˜ justin/

• Peter Røgen: Peter.Roegen*mat|dtu|dk,
• Dale Rolfsen: rolfsen*math|ubc|kuca,

http://www.math.ubc.ca/˜ rolfsen/

• Yongwu Rong:

• A.A.Rosly:

• Marc Rosso

• Colin Rourke: http://www.maths.warwick.ac.uk/˜ cpr/

• Lev Rozansky: http://www.math.unc.edu/Faculty/rozansky/ S: (see page 32)

• Masahico Saito: http://www.math.usf.edu/ saito/

• Joe Sawada: http://old-www.cis.uoguelph.ca/ sawada/

• Brian Sanderson: http://www.maths.warwick.ac.uk/˜ bjs/

• Uwe Sassenberg:

• Steve F. Sawin:

• Jörg Sawollek:

• Justin Sawon: http://www.math.sunysb.edu/˜ sawon/

• Kevin Patrick Scannell: scannell*slu|edu, http://borel.slu.edu/index.html

• Rob Schneiderman: http://www.math.upenn.edu/˜ schneiro/

• K.G.Selivanov:

• Karol Selwat: Karol|Selwat*math|uni|wroc|pl,

http://www.math.uni.wroc.pl/˜ kselwat/

• John Shareshian:

• Akiko Shima:

• Nadya Shirokova:

• Alexander Shumakovitch:

• Adam S. Sikora: http://www.math.buffalo.edu/Sikora Adam.html

• Dev Prakash Sinha: dps*math|uoregon|edu http://darkwing.uoregon.edu/˜ dps/ • Serguei A. Sirotine:

• Craig T. Snydal:

• E. Soboleva:

• Alexei Sossinsky: http://www.mccme.ru/˜ abs/

• Bill Spence:

• Ted Stanford:

• Ines Stassen:

• Alexander Stoimenow: http://www.ms.u-tokyo.ac.jp/˜ stoimeno/

• Yoshiaki Suetsugu

• Jacek Swiatkowski: http://www.math.uni.wroc.pl:80/˜ swiatkow/ T: (see page 34)

• Serge Tabachnikov:

• Akiko Tani:

• Kouki Taniyama: http://www.f.waseda.jp/taniyama/

• Clifford Henry Taubes:

• Peter Teichner: http://www.math.ucsd.edu/˜ teichner/

• George Thompson: http://www.ictp.trieste.it/ thompson/

• Dylan Paul Thurston: dpt*math|harvard|edu,

http://www.math.columbia.edu/ dpt/

• Victor Turchin (=Tourtchine): turchin*mccme|ru, http://www.math.ucl.ac.be/membres/turchin/

• Rolland Trapp:

• Tatsuya Tsukamoto: http://home.att.ne.jp/sun/tatsuya/

• Kyoichi Tsurusaki:

• Vladimir Turaev: http://www.cesj.com/turaev.html

• Svetlana D. Tyurina: See Svetlana Varchenko. U: (see page 35) V: (see page 35)

• Arkady Vaintrob: vaintrob*math|uoregon|edu,

http://darkwing.uoregon.edu/˜ vaintrob/

• Alexandre N. Varchenko: http://www.math.unc.edu/Faculty/av/ • Svetlana Varchenko: http://www.unc.edu/˜ svetlana/index.html • Victor A. Vassiliev: vva*mi|ras|ru

• Vladimir V. Vershinin:

• Oleg Viro: http://www.math.uu.se/˜ oleg/

• Pierre Vogel: http://www.math.jussieu.fr/˜ vogel

• Ismar Volic: http://www.people.virginia.edu/˜ iv2n/

• Alexander A. Voronov: http://www.math.umn.edu/˜ voronov/ W: (see page 37)

• Zhenghan Wang:

• Volkmar Welker: welker*mathematik|uni-marburg|de,

http://www.Mathematik.uni-marburg.de/˜ welker/

• Dominic J. A. Welsh:

• Simon Willerton: http://www.sheffield.ac.uk/simonwillerton/ • Edward Witten: http://www.sns.ias.edu/ witten/

• Jie Wu: http://www.math.nus.edu.sg/˜ matwujie/ X: (see page 37) Y: (see page 37)

• Harumi Yamada:

• Minoru Yamamoto: minomoto*kurume-nct|ac|jp
• Su-Win Yang: swyang*math|ntu|edu|tw

• Akira Yasuhara: http://www.u-gakugei.ac.jp/˜ yasuhara/

• David N. Yetter: http://www-personal.ksu.edu/ dyetter/ Z: (see page 38)

• Don Zagier:

• Jun Zhu:

• Gregg J. Zuckerman:

3. Acknowledgement

During the maintenance of this bibliography, comments and suggestions of many people were very helpful. In particular, we thank F. Aicardi, N. Askitas, G. Kuperberg, V. Liskovets, J.-B. Meilhan, D. Moskovich, K. Selwat, T. Stanford, A. Stoimenow, D. P. Thurston, A. Vaintrob, and S. Willerton.

4. References

4.1. References beginning with A.

(1) F. Aicardi, Topological invariants of knots and framed knots in the solid torus, C. R. Ac. Sci. Paris, Ser. 1, 321, no. 1, 81–86 (1995).

(2) F. Aicardi, Invariant polynomial of framed knots in the solid torus and its application to wave fronts and Legendrian knots. J. Knot Theory Ramifications 5 (1996), no. 6, 743–778.

(3) A. K. Aiston, Adams operators and knot decorations, q-alg/9711015 and Liverpool University preprint, November 1997.

(4) P. M. Akhmetiev, On a higher-order analog of the linking number of closed curves, September 1997.

(5) P. M. Akhmetiev and D. Repovs, A generalization of the Sato-Levine invariant, September 1997.

(6) S. Allyonov, Arrow diagram formulas for fourth order knot invariants, Fundamentalnaya i prikladnaya matematika, 2005, v. 11, no. 5 (2005), pp.3–17 (in Russian).

(7) S. Allyonov, On diagram formulas for knot invariants, Trans. Steklov Math. Inst., 2006, v. 252, pp. 10—17 (in Russian).

(8) D. Altschuler, Representations of knot groups and Vassiliev invariants, q-alg/9503015 preprint, March 1995.

(9) D. Altschuler and L. Freidel, On universal Vassiliev invariants, hep-th/9403053.

(10) D. Altschuler and L. Freidel, Vassiliev knot invariants and Chern-Simons perturbation theory to all orders, Comm. Math. Phys. 187 (1997) 261–287, arXiv:q-alg/9603010.

(11) M. Alvarez and J. M. F. Labastida, Numerical knot invariants of finite type from Chern-Simons perturbation theory, Nuclear Physics B 433 (1995) 555-596, arXiv:hep-th/9407076.

(12) M. Alvarez and J. M. F. Labastida, Vassiliev invariants for torus knots, Jour. of Knot Theory and its Ramifications 5(6) (1996) 779–803, arXiv:q-alg/9506009.

(13) M. Alvarez and J. M. F. Labastida, Primitive Vassiliev invariants and factorization in Chern-Simons perturbation theory, Comm. Math. Phys. 189-3 (1997), arXiv:q-alg/9604010.

(14) M. Alvarez, J. M. F. Labastida and E. Pérez, Vassiliev invariants for links from Chern-Simons perturbation theory, Nuclear Physics B 488 (1997) 677, arXiv:hep-th/9607030.

(15) J. E. Andersen and J. Mattes, Configuration space integrals and universal Vassiliev invariants over closed surfaces, q-alg/9704019 preprint, April 1997.

(16) J. E. Andersen, J. Mattes and N. Reshetikhin, The poisson structure on the moduli space of flat connections and chord diagrams, Topology 35 (1996) 1069–1083.

(17) J. E. Andersen, J. Mattes and N. Reshetikhin, Quantization of the algebra of chord diagrams, Mathematical Proceedings of the Cambridge Philosophical Society (1998), arXiv:q-alg/9701018

(18) J. E. Andersen and V. Turaev, Higher skein modules, Århus University and MaPhySto preprint, November 1998, arXiv:math.GT/9812071.

(19) E. Appleboim, Finite type invariants of links with fixed linking matrix, Jour. of Knot Theory and its Ramifications, to appear, arXiv:math.GT/9906138.

(20) V. I. Arnold, The Vassiliev theory of discriminants and knots, First European Congress of Mathematics I 3–29, Birkhäuser Basel 1994.

(21) N. A. Askitas, A formula in the theory of finite type invariants, University of Thessaloniki preprint, April 1999,arXiv:math.GT/9904141.

(22) N. A. Askitas and E. Kalfagianni, On knot adjacency, Topology and its Applications 126 (2002) 63–81.

(23) E. Auclair and C. Lescop, Clover calculus for homology 3-spheres via basic algebraic topology, Algebraic and Geometric Topology 5 (2005) 71–106, arXiv:math.GT/0401251.

(24) B. Audoux, Heegaard–Floer homology for singular knots, arXiv:0705.2377.

4.2. References beginning with B.

(1) S. Baader, A note on Vassiliev invariants of quasipositive knots, University of Basel preprint, December 2004, arXiv:math.GT/0412453.

(2) S. Baader, Gordian distance and Vassiliev invariants, arXiv:math/0703786.

(3) E. Babson, A. Björner, S. Linusson, J. Shareshian and V. Welker, Complexes of not i-connected graphs, Topology 38-2 (1999) 271–299, arXiv:math.CO/9705219.

(4) R. Bacher, Spin models for chord diagrams, Université de Grenoble I preprint, May 1998.

(5) J. C. Baez, Link invariants of finite type and perturbation theory, Lett. Math. Phys. 26 (1992) 43–51, arXiv:hep-th/9207041.

(6) D. Bar-Natan, Weights of Feynman diagrams and the Vassiliev knot invariants, Princeton University preprint, February 1991.

(7) D. Bar-Natan, On the Vassiliev knot invariants, Topology 34 (1995) 423–472.

(8) D. Bar-Natan, Vassiliev homotopy string link invariants, Jour. of Knot Theory and its Ramifications 4 (1995) 13–32.

(9) D. Bar-Natan, Non-associative tangles, in Geometric topology (proceedings of the Georgia international topology conference), (W. H. Kazez, ed.), 139–183, Amer. Math. Soc. and International Press, Providence, 1997.

(10) D. Bar-Natan, Some computations related to Vassiliev invariants, electronic publication (circa 1996), http://www.math.toronto.edu/˜ drorbn/LOP.html#Computations.

(11) D. Bar-Natan, Vassiliev and quantum invariants of braids, in Proc. of Symp. in Appl. Math. 51 (1996) 129–144, The interface of knots and physics, (L. H. Kauffman, ed.), Amer. Math. Soc., Providence.

(12) D. Bar-Natan, Polynomial invariants are polynomial, Math. Res. Lett. 2 (1995) 239–246. See also q-alg/9606025.

(13) D. Bar-Natan, Lie algebras and the Four Color Theorem, Combinatorica 17-1 (1997) 43–52, arXiv: q-alg/9606016.

(14) D. Bar-Natan, Bracelets and the Goussarov filtration of the space of knots, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 1–12, arXiv:math.GT/0111267.

(15) D. Bar-Natan, From astrology to topology via Feynman diagrams and Lie algebras, Rendiconti Del Circolo Matematico Di Palermo Serie II 63 (2000) 11–16, http://www.math.toronto.edu/

∼drorbn/Talks/Srni-9901.

(16) D. Bar-Natan, Finite type invariants, Toronto University preprint, August 2004, arXiv:math.GT/0408182.

(17) D. Bar-Natan and S. Garoufalidis, On the Melvin-Morton-Rozansky conjecture, Invent. Math. 125 (1996) 103–133.

(18) D. Bar-Natan, S. Garoufalidis, L. Rozansky and D. P. Thurston, Wheels, wheeling, and the Kontsevich integral of the unknot, Israel Journal of Mathematics 119 (2000) 217–237, arXiv:q-alg/9703025.

(19) D. Bar-Natan, S. Garoufalidis, L. Rozansky and D. P. Thurston, The Århus integral of rational homology 3-spheres I: A highly non trivial flat connection on S^{3}, Selecta Mathematica, New Series 8 (2002) 315–339, arXiv:q-alg/9706004.

(20) D. Bar-Natan, S. Garoufalidis, L. Rozansky and D. P. Thurston, The Århus integral of rational homology 3-spheres II: Invariance and universality, Selecta Mathematica, New Series 8 (2002) 341– 371, arXiv:math.QA/9801049.

(21) D. Bar-Natan, S. Garoufalidis, L. Rozansky and D. P. Thurston, The Århus integral of rational homology 3-spheres III: The relation with the Le-Murakami-Ohtsuki invariant, Selecta Mathematica, New Series 10 (2004) 305–324, arXiv:math.QA/9808013.

(22) D. Bar-Natan and R. Lawrence, A Rational Surgery Formula for the LMO Invariant, Israel Journal of Mathematics 140 (2004) 29–60, arXiv:math.GT/0007045.

(23) D. Bar-Natan, T. Q. T. Le and D. P. Thurston, Two applications of elementary knot theory to Lie algebras and Vassiliev invariants, Geometry and Topology 7-1 (2003) 1–31, arXiv:math.QA/0204311.

(24) D. Bar-Natan and A. Stoimenow, The fundamental theorem of Vassiliev invariants, in Proc. of the Århus Conf. Geometry and physics, (J. E. Andersen, J. Dupont, H. Pedersen, and A. Swann, eds.), lecture notes in pure and applied mathematics 184 (1997) 101–134, Marcel Dekker, New-York. See also arXiv:q-alg/9702009.

(25) P. Bellingeri and L. Funar, Braids on surfaces and finite type invariants, Université de Grenoble I and Université de Montpellier II preprint, October 2003, arXiv:math.GT/0309245.

(26) B. Berceanu and S. Papadima, Universal representations of braid and braid-permutation groups, arXiv:0708.0634.

(27) A.-B. Berger and I. Stassen, The skein relation for the (𝔤_{2}, V)-link invariant, Comm. Math. Helv. 75-1 (2000) 134–155, arXiv:math.QA/9806136.

(28) A.-B. Berger and I. Stassen, Skein relations for link invariants coming from exceptional Lie algebras, Bern University preprint, September 1998. See also math.QA/9901077.

(29) M. Berger, Hamiltonian dynamics generated by Vassiliev invariants, J. Physics A 34 (2001) 1363– 1374.

(30) M. Berger, Topological invariants in braid theory, Letters in Mathematical Physics 55 (2001) 181– 192.

(31) J. S. Birman, New points of view in knot theory, Bull. Amer. Math. Soc. 28 (1993) 253–287. See also math.GT/9304209.

(32) J. S. Birman, On the combinatorics of Vassiliev invariants, in Braid groups, knot theory and statistical mechanics II Adv. Ser. Math. Phys. 17 (C. N. Yang and M. L. Ge, eds.), World Scientific, New-Jersey 1994, 1–19.

(33) J. S. Birman, Vassiliev invariants of knots and links: a survey, transparencies and video of MSRI lecture, June 1998.

(34) J. S. Birman and X-S. Lin, Knot polynomials and Vassiliev’s invariants, Invent. Math. 111 (1993) 225–270.

(35) J. S. Birman and R. Trapp, Braided Chord Diagrams, Jour. of Knot Theory and its Ramifications 7(1) (1998) 1–22. See also math.GT/9804020.

(36) B. Bollobás and O. Riordan, Linearized chord diagrams and an upper bound for Vassiliev invariants, Jour. of Knot Theory and its Ramifications 9(7) (2000) 847–853.

(37) R. Bott, Configuration spaces and imbedding problems, in Proc. of the Århus Conf. Geometry and physics, (J. E. Andersen, J. Dupont, H. Pedersen, and A. Swann, eds.), lecture notes in pure and applied mathematics 184 (1997) 135–140, Marcel Dekker, New-York.

(38) R. Bott, Configuration space invariants of knots and 3-manifolds, transparencies and video of MSRI lecture, October 1998.

(39) R. Bott and C. Taubes, On the self-linking of knots, Jour. Math. Phys. 35 (1994).

(40) D. J. Broadhurst, Conjectured enumeration of Vassiliev invariants, Open University UK preprint, September 1997, arXiv:q-alg/9709031.

(41) D. J. Broadhurst and D. Kreimer, Feynman diagrams as a weight system: four-loop test of a fourterm relation, hep-th/9612011, Open University UK (OUT-4102-66), and Mainz University (MZ-TH/96-37) preprint, November 1996.

(42) R. Budney, J. Conant, K. P. Scannell and D. Sinha, New perspectives on self-linking, Advances in Mathematics, 191 (2005) 78–113. arXiv:math.GT/0303034.

(43) R. Budney. Topology of spaces of knots in dimension 3. arXiv:math.GT/0506524.

(44) D. Bullock, C. Frohman and J. Kania-Bartoszynska, Understanding the Kauffman bracket skein module, Jour. of Knot Theory and its Ramifications 8(3) (1999) 265–277, arXiv:q-alg/9604013.

(45) U. Burri, For a fixed Turaev shadow all “Jones-Vassiliev” invariants depend polynomially on the gleams, University of Basel preprint, March 1995.

(46) U. Burri, Vassiliev invariants and gleam polynomials, q-alg 9605019 preprint, May 1996.

4.3. References beginning with C.

(1) R. Campoamor-Stursberg and V. O. Manturov, Invariant Tensors Formulae via Chord Diagrams, Universidad Complutense and Moscow State University preprint, May 2005, arXiv:math.GT/0505320.

(2) P. Cartier, Construction combinatoire des invariants de Vassiliev-Kontsevich des nœuds, C. R. Acad. Sci. Paris 316 Série I (1993) 1205–1210.

(3) A. S. Cattaneo, Configuration space integrals and invariants for 3-manifolds and knots, Low Dimensional Topology, ed. H. Nencka, Cont. Math. 233, (1999) 153–165, arXiv:math.GT/9912083.

(4) A. S. Cattaneo, P. Cotta-Ramusino, J. Fröhlich and M. Martellini, Topological BF theories in 3 and 4 dimensions, J. Math. Phys. 36 (1995) 6137–6160, arXiv:hep-th/9505027.

(5) A. S. Cattaneo, P. Cotta-Ramusino and R. Longoni, Configuration spaces and Vassiliev classes in any dimension, Algebraic and Geometric Topology 2 (2002) 949–1000, arXiv:math.GT/9910139.

(6) A. S. Cattaneo, P. Cotta-Ramusino and R. Longoni, Algebraic structures on graph cohomology, Universität Zürich-Irchel, Università Degli Studi di Milano and Università di Roma “La Sapienza” preprint, July 2003, arXiv:math.GT/0307218.

(7) N. Chbili, On the invariants of lens knots, in Proc. Knots 96 (S. Suzuki, ed.) (1997) 365–375, World Scientific.

(8) V. Chernov, The most refined Vassiliev invariant of degree one of knots and links in ℝ^{1}-fibrations over a surface, Jour. of Knot Theory and its Ramifications 7(2) (1998) 257–266, arXiv:math.GT/9906137.

(9) V. Chernov, Arnold-type invariants of curves on surfaces, Jour. of Knot Theory and its Ramifications 8(1) (1999) 71–97, arXiv:math.GT/9906125.

(10) V. Chernov, Shadows of wave fronts and Arnold-Bennequin type invariants of fronts on surfaces and orbifolds, Amer. Math. Soc. Transl. 2-190 (1999) 159–184, arXiv:math.GT/9906121.

(11) V. Chernov, Finite order invariants of Legendrian, transverse, and framed knots in contact 3manifolds, ETH Zürich preprint, July 1999, arXiv:math.SG/9907118.

(12) V. Chernov, Vassiliev invariants of Legendrian, of transverse and framed knots in contact 3-manifolds, Max-Planck-Institut Bonn preprint, April 2000, arXiv:math.SG/0005002.

(13) V. Chernov, Isomorphism of the groups of Vassiliev invariants of Legendrian and of pseudo Legendrian Knots in contact 3-manifolds, Compositio Math., to appear, arXiv:math.GT/0012065.

(14) V. Chernov, The universal order one invariant of framed knots in most S^{1}-bundles over orientable surfaces, Algebraic and Geometric Topology 3-3 (2003) 89-101, arXiv:math.GT/0209027.

(15) S. V. Chmutov, A proof of the Melvin-Morton conjecture and Feynman diagrams, Jour. of Knot Theory and its Ramifications, 7(1) (1998) 23–40.

(16) S. V. Chmutov, Combinatorial analog of the Melvin-Morton conjecture, Proceedings of KNOTS ’96, World Scientific Publishing Co. (1997) 257–266.

(17) S. V. Chmutov and S. V. Duzhin, An upper bound for the number of Vassiliev knot invariants, Jour. of Knot Theory and its Ramifications, 3(2) (1994), 141–151.

(18) S. V. Chmutov and S. V. Duzhin, A lower bound for the number of Vassiliev knot invariants, Topology and Applications 92 (1999) 201–223.

(19) S. V. Chmutov and S. V. Duzhin, The Kontsevich integral, Acta Applicandae Mathematicae 66 (2001) 155–190.

(20) S. V. Chmutov and S. V. Duzhin, The Kontsevich integral, to appear in the Encyclopedia of Mathematical Physics, Elsevier, arXiv:math.GT/0501040.

(21) S. V. Chmutov, S. V. Duzhin and A. I. Kaishev, The algebra of 3-graphs, Trans. Steklov Math. Institute 221 (1998), 168–196.

(22) S. V. Chmutov, S. V. Duzhin and S. K. Lando, Vassiliev knot invariants I. Introduction, in Adv. in Soviet Math., 21 (1994) Singularities and curves, (V. I. Arnold, ed.), 117–126.

(23) S. V. Chmutov, S. V. Duzhin and S. K. Lando, Vassiliev knot invariants II. Intersection graph conjecture for trees, in Adv. in Soviet Math., 21 (1994) Singularities and curves, (V. I. Arnold, ed.), 127–134.

(24) S. V. Chmutov, S. V. Duzhin and S. K. Lando, Vassiliev knot invariants III. Forest algebra and weighted graphs, in Adv. in Soviet Math., 21 (1994) Singularities and curves, (V. I. Arnold, ed.), 135–145.

(25) S. Chmutov, S. Duzhin, J. Mostovoy. Introduction to Vassiliev Knot invariants, accepted for publication by Cambridge University Press, March 2011. Draft at http://arxiv.org/abs/1103.5628.

(26) S. V. Chmutov and V. Goryunov, Kauffman bracket of plane curves, Comm. Math. Physics 182 (1996) 83–103.

(27) S. V. Chmutov and A. N. Varchenko, Remarks on the Vassilliev knot invariants coming from sl_{2}, Topology 36-1 (1997).

(28) T. D. Cochran and P. M. Melvin, Finite type invariants of 3-manifolds, Rice University and Bryn Mawr College preprint, November 1997. See also math.GT/9805026.

(29) T. D. Cochran and P. M. Melvin, Quantum cyclotomic orders of 3-manifolds, Rice University and Bryn Mawr College preprint, November 1997. See also math.GT/9809129.

(30) F. .R. Cohen and J. Wu, Braid groups, free groups, and the loop space of the 2-sphere, University of Rochester and National University of Singapore preprint, September 2004, arXiv:math.AT/0409307.

(31) J. Conant, A knot bounding a grope of class n is ⌈\frac{n}{2}⌉-trivial, University of California at San Diego preprint, July 1999. See also math.GT/9907158.

(32) J. Conant, On a theorem of Goussarov, Jour. of Knot Theory and its Ramifications 12-1 (2003) 47–52, arXiv:math.GT/0110057.

(33) J. Conant, Vassiliev invariants and embedded gropes, Cornell University preprint, undated.

(34) J. Conant, Gropes and the rational lift of the Kontsevich integral, University of Tennessee preprint, April 2004, arXiv:math.GT/0404270.

(35) J. Conant, Chirality and the Conway polynomial, University of Tennessee preprint, March 2005, arXiv:math.GT/0503648.

(36) J. Conant, Homotopy approximations to the space of knots, Feynman diagrams, and a conjecture of Scannell and Sinha, arXiv:math.GT/0601647.

(37) J. Conant, R. Schneiderman and P. Teichner, Jacobi identities in low-dimensional topology, University of Tennessee, Courant Institute and University of California at San Diego preprint, January 2004, arXiv:math.GT/0401427.

(38) J. Conant and P. Teichner, Grope cobordism of classical knots, Topology 43-1) (2004) 119–156, arXiv:math.GT/0012118.

(39) J. Conant and P. Teichner, Grope cobordism and Feynman diagrams, Mathematische Annalen 328 (2004) 135–171, arXiv:math.GT/0209075.

(40) James Conant, Jacob Mostovoy and Ted Stanford, Finite type invariants based on the band-pass and doubled delta moves. arXiv:math.GT/0511189.

4.4. References beginning with D.

(1) O. T. Dasbach, On subspaces of the space of Vassiliev invariants, Düsseldorf University thesis, August 1995.

(2) O. T. Dasbach, A remark on the HOMFLY-Vassiliev invariants, Düsseldorf University preprint, October 1996.

(3) O. T. Dasbach, On the combinatorial structure of primitive Vassiliev invariants II, Jour. Comb. Theory, Ser. A, to appear.

(4) O. T. Dasbach, On the combinatorial structure of primitive Vassiliev invariants III — a lower bound, Comm. in Cont. Math. 2-4 (2000) 579–590, arXiv:math.GT/9806086.

(5) O. T. Dasbach and X-S. Lin, The Bennequin number of n-trivial closed n-braids is negative, University of California at Riverside preprint, October 2000, arXiv:math.GT/0010278.

(6) C. Day, Vassiliev invariants for links, Univ. of North Carolina at Chapel Hill preprint, 1992.

(7) J. Dean, Many classical knot invariants are not Vassiliev invariants, Jour. of Knot Theory and its Ramifications, 3(1) (1994) 7–9.

(8) C. De Concini and C. Procesi, Hyperplane arrangements and holonomy equations, Selecta Math., to appear.

(9) T. Deguchi and K. Tsurusaki, A statistical study of random knotting using the Vassiliev invariants, Jour. of Knot Theory and its Ramifications 3(3) (1994) 321–353.

(10) T. Deguchi and K. Tsurusaki, Numerical application of knot invariants and universality of random knotting, in Knot theory (V. F. R. Jones, J. Kania-Bartoszyńska, J. H. Przytycki, P. Traczyk, and V. G. Turaev, eds.), Banach Center Publications 42 77–85, Warsaw 1998.

(11) C. Di Bartolo, R. Gambini, J. Griego and J. Pullin, Consistent canonical quantization of general relativity in the space of Vassiliev knot invariants, gr-qc/9909063 preprint, September 1999.

(12) C. Di Bartolo, R. Gambini, J. Griego and J. Pullin, Canonical quantum gravity in the Vassiliev invariants arena: I. Kinematical structure, preprint, November 1999, arXiv:gr-qc/9911009.

(13) C. Di Bartolo, R. Gambini, J. Griego and J. Pullin, Canonical quantum gravity in the Vassiliev invariants arena: II. Constraints, habitats and consistency of the constraint algebra, preprint, November 1999, arXiv:gr-qc/9911010.

(14) Tudor Dimofte, Sergei Gukov, Jonatan Lenells, Don Zagier, Exact Results for Perturbative Chern-Simons Theory with Complex Gauge Group. http://arxiv.org/abs/0903.2472

(15) M. Domergue and P. Donato, Integrating a weight system of order n to an invariant of (n − 1)- singular knots, Jour. of Knot Theory and its Ramifications, 5(1) (1996) 23–35.

(16) S. Duzhin , Lectures on the Vassiliev knot invariants. Lectures in Mathematical Sciences, The University of Tokyo, vol. 19, 2002. 123 pp. http://www.pdmi.ras.ru/ duzhin/Vics/vics.ps.gz.

(17) S. Duzhin, Conway polynomial and Magnus expansion. http://arxiv.org/abs/1001.2500

(18) S. Duzhin and M.Karev, Detecting the orientation of long links by finite type invariants . Preprint, July 1, 2005. http://www.pdmi.ras.ru/ duzhin/papers/, arXiv:math.GT/0507015.

(19) I. Dynnikov, The Alexander polynomial in several variables can be expressed in terms of the Vassiliev invariants. Russian Mathematical Surveys, 1997, 52:1, 219–221.

4.5. References beginning with E.

(1) M. Eisermann, Knotengruppen-Darstellungen und Invarianten von endlichem Typ (Knot group representations and invariants of finite type), Ph.D. thesis, University of Bonn, January 2000.

(2) M. Eisermann, Les invariants rationnels de type fini ne distinguent pas les nœuds dans S^{2} × S^{1}, Comptes Rendus de l’Académie des Sciences de Paris, Série I 332 (2001) 51–55.

(3) M. Eisermann, The number of knot group representations is not a Vassiliev invariant, Proceedings of the American Mathematical Society 128 (2000) 1555–1561.

(4) M. Eisermann, A geometric characterization of Vassiliev invariants, ENS Lyon preprint, March 2001.

(5) M. Eisermann, Vassiliev invariants do not distinguish knots in a Whitehead manifold, ENS Lyon preprint, April 2001.

(6) M. Eisermann, Vassiliev invariants and the Poincaré Conjecture, ENS Lyon preprint, April 2001. Published version: Topology 43 (2004), 1211-1229 (MR 2080001).

(7) T. Ekholm, Invariants of generic immersions, Pac. Jour. of Math., to appear, arXiv:math.GT/9904031.

(8) T. Ekholm, Geometry of self intersection surfaces and Vassiliev invariants of generic immersions S^{k} → R^{2k−2}, Leningrad Math. Jour., to appear.

(9) T. Ekholm, Regular homotopy and Vassiliev invariants of generic immersions S^{k} → R^{2k−1}, Jour. of Knot Theory and its Ramifications, to appear.

(10) B. Enriquez and F. Gavarini, A formula for the logarithm of the KZ associator, IRMA Strasbourg and Università Degli Studi di Roma preprint, April 2004, arXiv:math.QA/0404262.

4.6. References beginning with F.

(1) R. Fenn, Vassiliev theory for knots, Tr. J. of Math. 18 (1994) 81–101.

(2) R. Fenn, D. Rolfsen and J. Zhu, Centralisers in the braid group and singular braid monoid, L’Enseignement Math. 42 (1996).

(3) R. Fenn, C. Rourke, and B. Sanderson, James bundles and applications, University of Sussex and University of Warwick preprint, January 1996.

(4) T. Fiedler, New invariants in knot theory, Université Toulouse III preprint no. 172, November 1999.

(5) T. Fiedler, Global knot theory in F^{2} × ℝ, preprint, December 2000, arXiv:math.GT/0012087.

(6) T. Fiedler, Gauss diagram invariants for knots which are not closed braids, Université Toulouse III preprint no. 216, May 2001.

(7) T. Fiedler, Gauss Diagram Invariants for Knots and Links, Kluwer Academic Publishers, 2001.

(8) T. Fiedler, One parameter knot theory. Université Toulouse preprint, November 2002. Revised: prepublication n. 262, Lab. de math., Universite Paul Sabatier, April 2003.

(9) T. Fiedler and A. Stoimenow, New knot and link invariants, Humboldt Berlin University preprint, December 1996.

(10) J. M. Figueroa-O’Farrill, T. Kimura and A. Vaintrob, The universal Vassiliev invariant for the Lie superalgebra gl(1 | 1), q-alg/9602014 preprint, February 1996.

(11) J. Fine, Vassiliev theory and regional change, math.QA/9803004 preprint, December 1996.

(12) V.V.Fock, N.A.Nekrasov, A.A.Rosly, and K.G.Selivanov (1992) What we think about the higherdimensional Chern-Simons theories, in Sakharov Memorial Lectures in Physics. V.1. Commack, NY: Nova Science Publishers, Inc., 465–471.

(13) L. Freidel, From sl(2) Kirby weight systems to the asymptotic 3-manifold invariant, Penn State University preprint, February 1998. See also math.QA/9802069

(14) L. Freidel and K. Krasnov, Spin Foam Models and the Classical Action Principle, Adv. Theor. Math. Phys. 2 (1999) 1183–1247, arXiv:hep-th/9807092.

(15) D. Fuchs and S. Tabachnikov, Invariants of Legendrian and transverse knots in the standard contact space, Topology 36-5 (1997) 1025–1053.

(16) L. Funar, Vassiliev invariants I: braid groups and rational homotopy theory, Revue Roumaine Math. Pures Appl., 42 (1997) 245–272, arXiv:q-alg/9510008.

(17) L. Funar, On knots having the same Vassiliev invariants up to a certain degree, Université de Grenoble I preprint, July 1999.

4.7. References beginning with G.

(1) R. Gambini, J. Griego and J. Pullin, A spin network generalization of the Jones polynomial and Vassiliev invariants, q-alg/9711014, University of Montevideo, and Pennsylvania State University preprint, November 1997.

(2) R. Gambini, J. Griego and J. Pullin, Vassiliev invariants: a new framework for quantum gravity, gr-qc/9803018, University of Montevideo, and Pennsylvania State University preprint, November (3) S. Garoufalidis, On finite type 3-manifold invariants I, Jour. of Knot Theory and its Ramifications

5 (1996) 441–462.

(4) S. Garoufalidis, A reappearance of wheels, Jour. of Knot Theory and its Ramifications 7 (1998) 1065–1071.

(5) S. Garoufalidis, Applications of the lantern identity, Brandeis University preprint, August 1997. See also q-alg/9708019.

(6) S. Garoufalidis, Signatures of links and finite type invariants of cyclic branched covers, to appear in the Rothenberg Festshift, Publications AMS. See also math.GT/9811021.

(7) S. Garoufalidis, The mystery of the brane relation, Georgia Institute of Technology preprint, October 1999, arXiv:math.GT/0006045.

(8) S. Garoufalidis, Whitehead doubling persists, Algebraic and Geometric Topology 4-40 (2004) 935– 942, arXiv:math.GT/0003189.

(9) S. Garoufalidis, Periodicity of Goussarov-Vassiliev knot invariants, Invariants of knots and 3manifolds (Kyoto 2001), Geometry and Topology Monographs 4 43-54, arXiv:math.GT/0201055.

(10) S. Garoufalidis, Beads: from Lie algebras to Lie groups, University of Warwick preprint, January 2002, arXiv:math.GT/0201056.

(11) S. Garoufalidis, Links with trivial Alexander module and nontrivial Milnor invariants, University of Warwick preprint, June 2002, arXiv:math.GT/0206196.

(12) S. Garoufalidis, Does the Jones polynomial determine the signature of a knot?, Georgia Institute of Technology preprint, October 2003, arXiv:math.GT/0310203.

(13) S. Garoufalidis, M. Goussarov and M. Polyak, Calculus of clovers and finite type invariants of 3-manifolds, Geometry and Topology 5-3 (2001) 75–108, arXiv:math.GT/0005192.

(14) S. Garoufalidis and N. Habegger, The Alexander polynomial and finite type 3-manifold invariants, to appear in Math. Annalen. See also q-alg/9708002.

(15) S. Garoufalidis and A. Kricker, A rational noncommutative invariant of boundary links, Geometry and Topology 8-4 (2004) 115–204, arXiv:math.GT/0105028.

(16) S. Garoufalidis and A. Kricker, Finite type invariants of cyclic branched covers, Georgia Institute of Technology preprint, July 2001, arXiv:math.GT/0107220.

(17) S. Garoufalidis and T. Q. T. Le, An analytic version of the Melvin-Morton-Rozansky Conjecture, Georgia Institute of Technology preprint, March 2005, arXiv:math.GT/0503641.

(18) S. Garoufalidis and J. Levine, On finite type 3-manifold invariants II, Math. Annalen 306 (1996) 691–718. See also q-alg/9506012.

(19) S. Garoufalidis and J. Levine, On finite type 3-manifold invariants IV: comparison of definitions, Math. Proc. Cambridge Phil. Soc. 122(2) (1997) 291–300. See also q-alg/9509028.

(20) S. Garoufalidis and J. Levine, Finite type 3-manifold invariants, the mapping class group and blinks, J. Diff. Geom. 47(2) (1997). See also q-alg/9712045.

(21) S. Garoufalidis and J. Levine, Finite type 3-manifold invariants and the structure of the Torelli group I, Invent. Math. 131(3) (1998) 541–594. See also q-alg/9603013.

(22) S. Garoufalidis and J. Levine, Tree-level invariants of three-manifolds, Massey products and the Johnson homomorphism, Harvard and Brandeis University preprint, April 1999. See also math.GT/9904106.

(23) S. Garoufalidis and J. Levine, Homology surgery and invariants of 3-manifolds, Georgia Institute of Technology and Brandeis University preprint, May 2000, arXiv:math.GT/0005280.

(24) S. Garoufalidis and J. Levine, Analytic invariants of boundary links, Georgia Institute of Technology and Brandeis University preprint, February 2001, arXiv:math.GT/0102103.

(25) S. Garoufalidis and J. Levine, Concordance and 1-loop clovers, Algebraic and Geometric Topology 1-33 (2001) 687–697, arXiv:math.GT/0102102.

(26) S. Garoufalidis and M. Loebl, Random walks and the colored Jones function, University of Warwick and Charles University preprint, March 2002, arXiv:math.GT/0203012.

(27) S. Garoufalidis and H. Nakamura, Some IHX-type relations on trivalent graphs and symplectic representation theory, Math. Res. Letters, 5 (1998) 391–402. See also q-alg/9705002.

(28) S. Garoufalidis and T. Ohtsuki, On finite type 3-manifold invariants III: manifold weight systems, Topology 37-2 (1998). See also q-alg/9705004.

(29) S. Garoufalidis and T. Ohtsuki, On finite type 3-manifold invariants V: rational homology 3-spheres, in Proc. of the Århus Conf. Geometry and physics, (J. E. Andersen, J. Dupont, H. Pedersen, and A. Swann, eds.), lecture notes in pure and applied mathematics 184 (1997) 445–457, Marcel Dekker, New-York.

(30) S. Garoufalidis and L. Rozansky, The loop expansion of the Kontsevich integral, the null move and S-equivalence, Georgia Institute of Technology and Yale University preprint, March 2000, arXiv: math.GT/0003187.

(31) S. Garoufalidis and P. Teichner, On Knots with trivial Alexander polynomial, Georgia Institute of Technology and University of California at San Diego preprint, June 2002, arXiv:math.GT/0206023.

(32) N. Geer, The Kontsevich integral and quantized Lie superalgebras, Georgia Institute of Technology preprint, November 2004, arXiv:math.GT/0411053.

(33) S. Gervais and N. Habegger, The Topological IHX Relation, Pure Braids, and the Torelli Group, Université de Nantes preprint, March 2000.

(34) J. González-Meneses and L. Paris, Vassiliev invariants for braids on surfaces, Universidad de Sevilla and Université de Bourgogne preprint, June 2000, arXiv:math.GT/0006014.

(35) V. Goryunov, Vassiliev invariants of knots in R^{3} and in a solid torus, in Differential and Symplectic Topology of Knots and Curves (S. Tabachnikov, ed.) AMS Translations Ser. 2 vol. 190 (1999), Providence.

(36) V. Goryunov, Vassiliev type invariants in Arnold’s J^{+}-theory of plane curves without direct selftangencies, to appear in Topology.

(37) V. Goryunov, Finite order invariants of framed knots in a solid torus and in Arnold’s J^{+}-theory of plane curves, in Proc. of the Århus Conf. Geometry and physics, (J. E. Andersen, J. Dupont, H. Pedersen, and A. Swann, eds.), lecture notes in pure and applied mathematics 184 (1997) 549– 556, Marcel Dekker, New-York.

(38) M. Goussarov, A new form of the Conway-Jones polynomial of oriented links, Zapiski nauch. sem. POMI 193 (1991) 4–9 (English translation in Topology of manifolds and varieties (O. Viro, editor), Amer. Math. Soc., Providence 1994, 167–172).

(39) M. Goussarov, On n-equivalence of knots and invariants of finite degree, Zapiski nauch. sem. POMI 208 (1993) 152–173 (English translation in Topology of manifolds and varieties (O. Viro, editor), Amer. Math. Soc., Providence 1994, 173–192).

(40) M. Goussarov, Interdependent modifications of links and invariants of finite degree, Topology 37-3 (1998) 595–602.

(41) M. Goussarov, Finite type invariants are presented by Gauss diagram formulas, Sankt-Petersburg Department of Steklov Mathematical Institute preprint (translated from Russian by O. Viro), December 1998.

(42) M. Goussarov, Finite type invariants and n-equivalence of 3-manifolds, C. R. Acad. Sci. Paris Sér. I Math. 329(6) (1999) 517–522.

(43) M. Goussarov, Variations of knotted graphs, geometric technique of n-equivalence, St. Petersburg Math. J. 12-4 (2001).

(44) M. Goussarov, M. Polyak and O. Viro, Finite type invariants of classical and virtual knots, Topology 39 (2000) 1045–1068, arXiv:math.GT/9810073.

(45) M. Graña and V. Turaev, Knot theory for self-indexed graphs, University of Strasbourg and Universidad de Buenos Aires preprint, April 2003, arXiv:math.GT/0304061.

(46) M. Greenwood and X-S. Lin, On Vassiliev knot invariants induced from finite type 3-manifold invariants Trans. of Amer. Math. Soc. 351-9 (1999) 3659–3672, arXiv:q-alg/9506001.

(47) Grishanov S.A., V.Meshkov, V.A.Vassiliev. Recognizing textile structures by finite type knot invariants, Journal of knot theory and its ramifications, 2009. v. 18 (2), p. 209-235.

(48) Grishanov S.A., V.A. Vassiliev. Two constructions of weight systems from invariants of knots in non-trivial 3-manifolds, Topology and its Applications, 2008, Volume 150, Issue 16, Pages 1757-1765.

(49) Grishanov S.A., V.A. Vassiliev. Fiedler type combinatorial formulas for generalized Fiedler type invariants of knots in M^{2} × R^{1}, Topology and its Applications, 2009, vol. 156, p. 2307-2316.

(50) Grishanov S.A., V.A. Vassiliev. Invariants of links in 3-manifolds and splitting problem of textile structures http://www.pdmi.ras.ru/˜ arnsem/papers/4gv.pdf to appear in Journal of Knot Theory and its Ramifications/2011/02.

(51) Grishanov S.A., V.A. Vassiliev. Gauss Diagram Invariants of Links in M^{2}×R^{1} http://www.pdmi.ras.ru/˜ arnsem/papers/5gv.pdf, to appear in Journal of Knot Theory and its Ramifications/2011/02.

4.8. References beginning with H.

(1) N. Habegger, Finite type 3-manifold invariants: a proof of a conjecture of Garoufalidis, Université de Nantes preprint, July 1996.

(2) N. Habegger, A computation of the universal quantum 3-manifold invariant for manifolds of rank greater than 2, Université de Nantes preprint, December 1996.

(3) N. Habegger, The topological IHX relation, Université de Nantes preprint, May 1998.

(4) N. Habegger, Milnor, Johnson, and Tree Level Perturbative Invariants, Université de Nantes preprint, March 2000.

(5) N. Habegger and A. Beliakova, The Casson-Walker-Lescop invariant as a quantum 3-manifold invariant, Université de Nantes and University of Berne preprint, August 1997. See also q-alg/9708029.

(6) N. Habegger and G. Masbaum, The Kontsevich integral and Milnor’s invariants, Université de Nantes and Université Paris 7 preprint, January 1997. Published in: Topology, vol. 39, pp. 1253– 1289 (2000).

(7) N. Habegger and K. E. Orr, Milnor link invariants and quantum 3-manifold invariants, Comm. Math. Helv. 74-2 (1999) 322-344.

(8) N. Habegger and G. Thompson, The universal perturbative quantum 3-manifold invariant, Rozansky-Witten invariants, and the generalized Casson invariant, Université de Nantesand ICTP (Trieste) preprint, November 1999. See also math.GT/9911049.

(9) K. Habiro, Claspers and finite type invariants of links, Geometry and Topology 4 (2000) 1–83.

(10) K. Habiro, Brunnian links, claspers and Goussarov-Vassiliev finite type invariants. arXiv:math.GT/0510460.

(11) K. Habiro, T. Kanenobu, and A. Shima, Finite type invariants of ribbon 2-knots, Low Dimensional Topology, ed. H. Nencka, Cont. Math. 233, (1999) 187–196.

(12) K. Habiro, J.-B. Meilhan, Finite type invariants and Milnor invariants for Brunnian links. arXiv:

math.GT/0510534.

(13) K. Habiro, J.-B. Meilhan, On the Kontsevich integral of Brunnian links. arXiv:math/0605312.

(14) K. Habiro and A. Shima, Finite type invariants of ribbon 2-knots, II, University of Tokyo preprint, November 1998.

(15) A. Haviv, Towards a diagrammatic analogue of the Reshetikhin-Turaev link invariants, Hebrew University PhD thesis, September 2002, arXiv:math.QA/0211031.

(16) N. Hayashi, Graph invariants of Vassiliev type and applications to 4D quantum gravity, q-alg/9503010 preprint, March 1995.

(17) L. Helme-Guizon, The coefficients of the HOMFLYPT and the Kauffman polynomials are pointwise limits of Vassiliev invariants, George Washington University preprint, March 2005, arXiv: math.GT/0503185.

(18) V. Hinich and A. Vaintrob, Cyclic operads and algebra of chord diagrams, Selecta Mathematica, New Series, 8 (2002) 237–282, arXiv:math.QA/0005197.

(19) A. C. Hirshfeld and U. Sassenberg, Derivation of the total twist from Chern-Simons theory, Jour.

of Knot Theory and its Ramifications 5(6) (1996) 805–847.

(20) A. C. Hirshfeld and U. Sassenberg, Explicit formulation of a third order finite knot invariant derived from Chen-Simons theory, Universität Dortmund preprint, November 1995.

(21) A. C. Hirshfeld, U. Sassenberg and T. Klöker, A combinatorial link invariant of finite type derived from Chern-Simons theory, Jour. of Knot Theory and its Ramifications 6(2) (1997) 243–280.

(22) H. Howards and J. Luecke, Strongly n-trivial knots, Wake Forest University and University of Texas preprint, April 2000. See also math.GT/0004183.

(23) Y. Huh and G. T. Jin, θ-curve polynomials and finite-type invariants, SAIT and KAIST preprint, April 2001, arXiv:math.GT/0104179.

(24) M. Hutchings, Integration of singular braid invariants and graph cohomology, Transactions of the AMS 350 (1998) 1791–1809.

4.9. References beginning with I.

4.10. References beginning with J.

(1) T. Januszkiewicz and J. Swiatkowski Finite type invariants of generic immersions of M^{n} into R^{2n} are trivial, in Differential and Symplectic Topology of Knots and Curves (S. Tabachnikov, ed.) AMS Translations Ser. 2 vol. 190 (1999), Providence.

(2) M.-J. Jeong and C.-Y. Park, Vassiliev invariants and double dating tangles, to appear in Jour. of Knot Theory and its Ramifications.

(3) M.-J. Jeong and C.-Y. Park, Polynomial invariants and Vassiliev invariants, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 89–101, arXiv:math.GT/0211045.

(4) G. T. Jin, Finite type invariants of knots, KAIST preprint (in Korean).

(5) G. T. Jin and J. H. Lee, Coefficients of HOMFLY polynomial and Kauffman polynomial are not finite type invariants, KAIST preprint, October 2001.

(6) J. Johannes, A type 2 polynomial invariant of links derived from the Casson-Walker invariant, Indiana University preprint, October 1997.

(7) P. M. Johnson, Vassiliev invariants of HOMFLY and Kauffman type, to appear in Jour. of Knot Theory and its Ramifications.

4.11. References beginning with K.

(1) E. Kalfagianni, Finite type invariants for knots in three manifolds, Topology 37-3 (1998) 673–707.

(2) E. Kalfagianni, Homology spheres with the same finite type invariants of bounded orders, Math. Res. Lett. 4 (1997) 341–347.

(3) E. Kalfagianni, Vassiliev invariants and orientation of pretzel knots, Jour. of Knot Theory and its Ramifications 7(2) (1998) 173–185.

(4) E. Kalfagianni, Surgery n-triviality and companion tori, Jour. of Knot Theory and its Ramifications 13-4 (2004) 441–456.

(5) E. Kalfagianni, Alexander polynomial, finite type invariants and volume of hyperbolic knots, Algebraic and Geometric Topology 4-48 (2004) 1111–1123, arXiv:math.GT/0411384.

(6) E. Kalfagianni and X-S. Lin, The HOMFLY polynomial for links in rational homology 3-spheres, Topology 38-1 (1999) 95–115, arXiv:q-alg/9509010.

(7) E. Kalfagianni and X-S. Lin, Regular Seifert surfaces and Vassiliev knot invariants, Rutgers University and University of California at Riverside preprint, June 1997, arXiv:math.GT/9804032.

(8) E. Kalfagianni and X-S. Lin, Milnor and finite type invariants of plat-closures, Math. Res. Lett. 5 (1998) 293–304, arXiv:math.GT/9804030.

(9) E. Kalfagianni and X-S. Lin, Knot adjacency and satellites, Topology and its Applications 138 (2004) 207–217, arXiv:math.GT/0308168.

(10) E. Kalfagianni and X-S. Lin (with an appendix by D. McCullough), Knot adjacency, genus and essential tori, Michigan State University and University of California at Riverside preprint, March 2004, arXiv:math.GT/0403024.

(11) E. Kalfagianni and X-S. Lin, Higher degree knot adjacency as obstruction to fibering, Michigan State University and University of California at Riverside preprint, March 2004, arXiv:math.GT/0403026.

(12) E. Kalfagianni and X-S. Lin, Seifert surfaces, Commutators and Vassiliev invariants, To appear in the J. of Knot Theory and its Ramifications (special volume in honor of L. Kauffman). arXiv: 0709.1689.

(13) T. Kanenobu, Vassiliev-type invariants of a theta-curve, Jour. of Knot Theory and its Ramifications 6(4) (1997) 455-477.

(14) T. Kanenobu, Kauffman polynomials as Vassiliev link invariants, in Proc. Knots 96 (S. Suzuki, ed.) (1997) 411–431, World Scientific.

(15) T. Kanenobu, Vassiliev knot invariants of order 6, Osaka City University preprint, June 1998.

(16) T. Kanenobu and Y. Miyazawa, HOMFLY polynomials as Vassiliev link invariants, in Knot theory (V. F. R. Jones, J. Kania-Bartoszyńska, J. H. Przytycki, P. Traczyk, and V. G. Turaev, eds.), Banach Center Publications 42 165–185, Warsaw 1998.

(17) T. Kanenobu and Y. Miyazawa and A. Tani, Vassiliev link invariants of order three, Jour. of Knot Theory and its Ramifications 7(4) (1998) 433–462.

(18) M. Kapranov, Rozansky-Witten invariants via Atiyah classes, Compositio Math. 115 (1999) 71–113, arXiv:alg-geom/9704009.

(19) C. Kassel, Quantum groups, Springer-Verlag GTM 155, Heidelberg 1994.

(20) C. Kassel, M. Rosso and V. Turaev, Quantum groups and knot invariants, Panoramas et synthèses 5, Société Mathématique de France, 1997.

(21) C. Kassel and V. Turaev, Chord diagram invariants of tangles and graphs, Duke Math. J. 92 (1998) 497–552.

(22) L. H. Kauffman, Knots and physics (second edition), World Scientific, Singapore 1993.

(23) L. H. Kauffman, Vassiliev invariants and the Jones polynomial, Quantum groups, integrable statistical models and knot theory (Nankai) (M. L. Ge and H. J. de Vega, eds.), World Scientific, Singapore 1993.

(24) L. H. Kauffman, Vassiliev invariants and the loop states in quantum gravity, Proceedings of conference on quantum gravity (Riverside) (J. Baez, ed.), Oxford University press, Oxford 1994.

(25) L. H. Kauffman, Functional integration and the Kontsevich integral, University of Illinois at Chicago preprint, April 1998. See also math.GT/9811137.

(26) L. H. Kauffman, Virtual knot theory, University of Illinois at Chicago preprint, November 1998. See also math.GT/9811028.

(27) L. H. Kauffman, Knot diagrammatics, University of Illinois at Chicago preprint, October 2004, arXiv:math.GT/0410329.

(28) L. H. Kauffman and Y. B. Magarshak, Vassiliev knot invariants and the structure of RNA folding, Knots and applications (L. H. Kauffman, ed.), World Scientific, Singapore 1994.

(29) L. H. Kauffman and Y. B. Magarshak, Graph invariants and the topology of RNA folding, Jour. of Knot Theory and its Ramifications 3(3) (1994) 233–245.

(30) L. H. Kauffman, M. Saito and S. F. Sawin, On finiteness of certain Vassiliev invariants, Jour. of Knot Theory and its Ramifications 6(2) (1997) 291–297. See also q-alg/9507035.

(31) A. Kawauchi, A survey of knot theory, Birkhäuser, Basel 1996.

(32) M.Kazarian, First order invariants of the strangeness type for plane curves, Transact. (Trudy) Steklov Math. Inst. 221 (1998), 213–224.

(33) M. Khovanov, Surfaces in three-space and braided algebras, Yale University preprint, April 1995.

(34) T. Kimura, A. A. Voronov and G. J. Zuckerman, Homotopy Gerstenhaber algebras and topological field theory, q-alg/9602009 preprint, February 1996.

(35) P. Kirk and C. Livingston, Vassiliev invariants of two component links and the Casson-Walker invariant, Topology 36-6 (1997) 1333–1353.

(36) P. Kirk and C. Livingston, Type 1 invariants in irreducible 3-manifolds, Indiana University preprint, August 1996. Pac. Jour. of Math. 183-2 (1998) 305–331.

(37) P. Kirk, C. Livingston and Z. Wang, The Gassner representation for string links, Indiana University preprint, April 1998. See also math.GT/9806035.

(38) J. A. Kneissler, The number of primitive Vassiliev invariants up to degree twelve, University of Bonn preprint, June 1997, arXiv:q-alg/9706022.

(39) J. A. Kneissler, Relations in the algebra Λ, University of Bonn preprint, October 1997.

(40) J. A. Kneissler, On spaces of connected graphs I: Properties of ladders, Proc. Inter. Conf. Knots in Hellas ’98, Series on Knots and Everything 24 (2000) 252–273, arXiv:math.QA/0301018.

(41) J. A. Kneissler, On spaces of connected graphs II: Relations in the algebra Λ, Joura. of Knot Theory and its Ramifications 10-5 (2001) 667–674, arXiv:math.QA/0301019.

(42) J. A. Kneissler, On spaces of connected graphs III: The ladder filtration, Joura. of Knot Theory and its Ramifications 10-5 (2001) 675–686, arXiv:math.QA/0301020.

(43) J. A. Kneissler, Die kombinatorik der diagrammalgebren von invarianten endlichen typs, Ph.D. thesis, Rheinische Friedrich-Wilhelms-Universität Bonn.

(44) I. Kofman and X-S. Lin, Vassiliev invariants and the cubical knot complex, Topology 42 (2003) 83–101, arXiv:math.GT/0010009.

(45) I. Kofman and Y. Rong, Approximating Jones coefficients and other link invariants by Vassiliev invariants, University of Maryland and George Washington University preprint, June 1999, arXiv:math.GT/0007093.

(46) T. Kohno, Vassiliev invariants and de-Rham complex on the space of knots, Contemp. Math. 179 (1994) 123–138.

(47) T. Kohno, Elliptic KZ system, braid group of the torus and Vassiliev invariants, University of Tokyo preprint, January 1996.

(48) T. Kohno, Loop spaces of configuration spaces and finite type invariants, Invariants of knots and 3manifolds (Kyoto 2001), Geometry and Topology Monographs 4 143–160, arXiv:math.AT/0211056.

(49) T. Kohno, Conformal field theory and topology, Iwanami Series in Modern Mathematics, Translations of Mathematical Monographs 210, Amer. Math. Soc., Providence 2002.

(50) M. Kontsevich, Vassiliev’s knot invariants, Adv. in Sov. Math., 16(2) (1993) 137–150.

(51) M. Kontsevich, Feynman diagrams and low-dimensional topology, First European Congress of Mathematics II 97–121, Birkhäuser Basel 1994.

(52) M. Kontsevich, Rozansky-Witten invariants via formal geometry, Compositio Math. 115 (1999) 115–127, arXiv:dg-ga/9704009.

(53) M.Kontsevich, Formal (non-)commutative symplectic geometry. In: L. Corvin, I. Gel’fand, J. Lepovsky (eds.), The I. M. Gel’fand’s mathematical seminars 1990–1992, Birkhäuser, Basel, 173–187 (1993).

(54) D. Kreimer, Weight systems from Feynman diagrams, Jour. of Knot Theory and its Ramifications 7(1) (1998) 61–85. See also hep-th/9612010.

(55) D. Kreimer, Knots and Feynman diagrams, Cambridge Lecture Notes in Physics 13, Cambridge University Press, Cambridge 2001.

(56) A. Kricker, Alexander-Conway limits of many Vassiliev weight systems, Jour. of Knot Theory and its Ramifications 6(5) (1997) 687–714. See also q-alg/9603011.

(57) A. Kricker, Covering spaces over claspered knots, Tokyo Institute of Technology preprint, December 1998. See also math.GT/9901029.

(58) A. Kricker, Branched cyclic covers and finite type invariants, Tokyo Institute of Technology preprint, January 2000. See also math.GT/0003035.

(59) A. Kricker, The lines of the Kontsevich integral and Rozansky’s rationality conjecture, Tokyo Institute of Technology preprint, May 2000, arXiv:math.GT/0005284.

(60) A. Kricker, A surgery formula for the 2-loop piece of the LMO invariant of a pair, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 161-181, arXiv: math.GT/0211057.

(61) A. Kricker and B. Spence, Ohtsuki’s invariants are of finite type, Jour. of Knot Theory and its Ramifications 6(4) (1997) 583–597. See also q-alg/9608007.

(62) A. Kricker, B. Spence and I. Aitchison, Cabling the Vassiliev invariants, Jour. of Knot Theory and its Ramifications 6(3) (1997) 327–358. See also q-alg/9511024.

(63) G. Kuperberg, Detecting knot invertibility, Jour. of Knot Theory and its Ramifications 5 (1996) 173–181. See also q-alg/9712048.

(64) G. Kuperberg and D. P. Thurston, Perturbative 3-manifold invariants by cut-and-paste topology, University of California at Davis and University of California at Berkeley preprint, December 1999, arXiv:math.GT/9912167.

(65) V. Kurlin, Compressed Drinfeld associators, University of Burgundy preprint, August 2004, arXiv: math.GT/0408398, Journal of Algebra 292 (2005), no. 1, 184-242.

4.12. References beginning with L.

(1) J. M. F. Labastida, Topological quantum field theory: a progress report, hep-th/9511037 preprint, November 1995.

(2) J. M. F. Labastida, Chern-Simons Gauge Theory: Ten Years After, Santiago preprint, US-FT-7-99, April 1999. See also hep-th/9905057.

(3) J. M. F. Labastida, Knot theory from a Chern-Simons gauge theory point of view, Santiago preprint, US-FT-3/00, February 2000. See also hep-th/0002221.

(4) J. M. F. Labastida, Knot invariants and Chern-Simons theory, Santiago preprint, July 2000, arXiv:hep-th/0007152.

(5) J. M. F. Labastida and E. Pérez, Kontsevich integral for Vassiliev invariants from Chern-Simons perturbation theory in the light-cone gauge, Journal of Mathematical Physics 39 (1998) 5183–5198. See also hep-th/9710176.

(6) J. M. F. Labastida and E. Pérez, Gauge-Invariant Operators for Singular Knots in Chern-Simons Gauge Theory, Nuclear Physics B, 527 (1998) 499. See also hep-th/9712139.

(7) J. M. F. Labastida and E. Pérez, Combinatorial formulae for Vassiliev invariants from Chern-Simons gauge theory, CERN preprint CERN-TH/98-193 and Universidade de Santiago de Compostela preprint US-FT-11/98. See also hep-th/9807155.

(8) J. M. F. Labastida and E. Pérez, Vassiliev invariants in the context of Chern-Simons gauge theory, Universidade de Santiago de Compostela preprint US-FT-18/98. See also hep-th/9812105.

(9) Pascal Lambrechts, Victor Tourtchine, Ismar Volic. The rational homology of spaces of long knots in codimension ¿2, arXiv:math.AT/0703649 preprint, March 2007.

(10) S. K. Lando, On primitive elements in the bialgebra of chord diagrams, Max-Planck-Institut Bonn preprint 94-47, September 1995.

(11) S. K. Lando, On a Hopf algebra in graph theory,. J. of Comb. Theory, Series B, 80 (2000), no. 1, 104–121.

(12) Lando, S. J-invariants of ornaments and decorated chord diagrams, Funct. Anal. and its Appl., to appear

(13) J. Lannes, Sur les invariants de Vassiliev de degré inférieur ou égal à 3, L’Ens. Math. 39 (1993), 295–316.

(14) T. Q. T. Le, An invariant of integral homology 3-spheres which is universal for all finite type invariants, in Solitons, geometry and topology: on the crossroad, (V. Buchstaber and S. Novikov, eds.) AMS Translations Series 2, Providence, arXiv:q-alg/9601002.

(15) T. Q. T. Le, On denominators of the Kontsevich integral and the universal perturbative invariant of 3-manifolds, Inv. Math. 135-3 (1999) 689–722, arXiv:q-alg/9704017.

(16) T. Q. T. Le, On perturbative P SU(n) invariants of rational homology 3-spheres, arXiv:math.GT/9802032 preprint, February 1998.

(17) T. Q. T. Le, The Lê-Murakami-Ohtsuki invariant, SUNY at Buffalo preprint, June 1999.

(18) T. Q. T. Le, Finite type invariants of 3-manifolds. arXiv:math.GT/0507145.

(19) T. Q. T. Le, H. Murakami, J. Murakami and T. Ohtsuki, A three-manifold invariant derived from the universal Vassiliev-Kontsevich invariant, Proc. Japan Acad. 71 Ser. A (1995) 125–127.

(20) T. Q. T. Le, H. Murakami, J. Murakami and T. Ohtsuki, A three-manifold invariant via the Kontsevich integral, Max-Planck-Institut Bonn preprint, 1995.

(21) T. Q. T. Le and J. Murakami, On Kontsevich’s integral for the HOMFLY polynomial and relations of multiple zeta numbers, Topology and its Applications 62 (1995) 193–206.

(22) T. Q. T. Le and J. Murakami, Kontsevich integral for Kauffman polynomial, Max-Planck-Institut Bonn preprint 93-33, 1993.

(23) T. Q. T. Le and J. Murakami, Representation of the category of tangles by Kontsevich’s iterated integral, Comm. Math. Phys. 168 (1995) 535–562.

(24) T. Q. T. Le and J. Murakami, The universal Vassiliev-Kontsevich invariant for framed oriented links, Compositio Math. 102 (1996), 41–64, arXiv:hep-th/9401016.

(25) T. Q. T. Le and J. Murakami, Parallel version of the universal Vassiliev-Kontsevich invariant, J. Pure and Appl. Algebra 121 (1997) 271–291.

(26) T. Q. T. Le, J. Murakami and T. Ohtsuki, On a universal quantum invariant of 3-manifolds, Topology 37-3 (1998) 539–574, arXiv:q-alg/9512002.

(27) C. Lescop, Introduction to the Kontsevich integral of framed tangles, CNRS Institut Fourier preprint, June 1999.

(28) C. Lescop, About the uniqueness and the denominators of the Kontsevich Integral, CNRS Institut Fourier preprint, April 2000. See also math.GT/0004094.

(29) C. Lescop, On configuration space integrals for links, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 183-199, arXiv:math.GT/0211062.

(30) C. Lescop, On the Kontsevich-Kuperberg-Thurston construction of a configuration-space invariant for rational homology 3-spheres, CNRS Institut Fourier preprint, October 2004, arXiv:math.GT/0411088.

(31) C. Lescop, Splitting formulae for the Kontsevich-Kuperberg-Thurston invariant for rational homology 3-spheres, CNRS Institut Fourier preprint, October 2004, arXiv:math.GT/0411431.

(32) J. Levine, Pure braids, a new subgroup of the mapping class group and finite-type invariants, Brandeis University preprint, September 1998. See also math.GT/9712221.

(33) J. Levine, Homology cylinders: an expansion of the mapping class group, Brandeis University preprint, October 2000, arXiv:math.GT/0010247.

(34) B-H. Li and H-W. Sun, Exact number of chord diagrams and an estimation of the number of spine diagrams of order n, Ch. Sci. Bull., 42-9 (1997) 705–718.

(35) J. Lieberum, Chromatic weight systems and the corresponding knot invariants, Math. Ann. 317-3 (2000) 459–482, arXiv:q-alg/9701013.

(36) J. Lieberum, On Vassiliev invariants not coming from semisimple Lie algebras, Jour. of Knot Theory and its Ramifications 8-5 (1999) 659–666, arXiv:q-alg/9706005.

(37) J. Lieberum, The number of independent Vassiliev invariants in the Homfly and Kauffman polynomials, Documenta Mathematica 5 (2000) 275–299, arXiv:math.QA/9806064.

(38) J. Lieberum, Invariants de Vassiliev pour les entrelacs dans S^{3} et dans les variétés de dimension trois, Ph.D. thesis, University of Strasbourg, September 1998.

(39) J. Lieberum, A skein module of links in cylinders over surfaces, Int. Jour. of Math. and Math. Sci. 32-9 (2002) 515–554, arXiv:math.QA/9911174.

(40) J. Lieberum, The LMO-invariant of 3-manifolds of rank one and the Alexander polynomial, Math. Ann. 318-4 (2000) 761–776, arXiv:math.QA/0002040.

(41) J. Lieberum, Universal Vassiliev invariants of links in coverings of 3-manifolds, University of Basel preprint, April 2001, arXiv:math.QA/0105019.

(42) J. Lieberum, The Drinfeld associator of gl(1|1), University of Basel preprint, April 2002, arXiv: math.QA/0204346.

(43) X-S. Lin, Vertex models, quantum groups and Vassiliev’s knot invariants, Columbia University preprint, 1991.

(44) X-S. Lin, Finite type invariants of 3-manifolds, Topology 33-1 (1994) 45–71.

(45) X-S. Lin, Milnor link invariants are all of finite type, Columbia University preprint, 1992.

(46) X-S. Lin, Power series expansions and invariants of links, in Geometric topology (proceedings of the Georgia international topology conference), (W. H. Kazez, ed.), 184–202, Amer. Math. Soc. and International Press, Providence, 1997.

(47) X-S. Lin, Knot invariants and iterated integrals, Columbia University preprint, 1994.

(48) X-S. Lin, Braid algebras, trace modules and Vassiliev invariants, Columbia University preprint, 1994.

(49) X-S. Lin, Invariants of Legendrian knots, Columbia University preprint, 1995.

(50) X-S. Lin, Finite type invariants of integral homology spheres: a survey, in Knot theory (V. F. R. Jones, J. Kania-Bartoszyńska, J. H. Przytycki, P. Traczyk, and V. G. Turaev, eds.), Banach Center Publications 42 205–220, Warsaw 1998. See also q-alg/9510003.

(51) X-S. Lin, Finite type link invariants and the non-invertibility of links, arXiv:q-alg/9601019. Published in Math. Res. Lett. 3-3 (1996) 405–417 under the title Finite type link invariants and the invertibility of links.

(52) X-S. Lin, Knot energies and knot invariants, Chaos, Solitons and Fractals 9 (1998) 645–655, arXiv:qalg/9604018.

(53) X-S. Lin, Melvin-Morton conjecture: a survey, Proceedings of ICCM 98, to appear.

(54) X-S. Lin, Link-homotopy invariants of finite type, University of California at Riverside preprint, December 2000, arXiv:math.GT/0012095. Published in l’Enseignement Mathematique, 47 (2001), pp. 315–327 under the title Finite type link-homotopy invariants.

(55) X-S. Lin and Z. Wang, Integral geometry of plane curves and knot invariants, Jour. of Diff. Geom. 44-1 (1996) 74–95, arXiv:dg-ga/9411015.

(56) X-S. Lin and Z. Wang, On Ohtsuki’s invariants of integral homology 3-spheres, I, Acta Math. Sinica (Ser. B) 15-3 (1999) 293–316, arXiv:q-alg/9509009.

(57) X-S. Lin and Z. Wang, Fermat limit and congruence of Ohtsuki invariants, Proceedings of the Kirbyfest, Geometry and Topology Monographs 2 321–333, arXiv:math.GT/9810147.

(58) R. Longoni, Sviluppo perturbativo delle teorie di campo topologiche di tipo BF e invarianti dei nodi, Master’s thesis (in Italian).

4.13. References beginning with M.

(1) S. A. Major, Embedded graph invariants in Chern-Simons theory, Universität Wien and Deep Springs College preprint, October 1998. See also hepth/9810071.

(2) V. O. Manturov, Vassiliev invariants for virtual links, curves on surfaces and the Jones-Kauffman polynomial, Moscow State Univ. preprint, July 2003.

(3) V. O. Manturov, Teoriya uzlov, (”Knot theory”, in Russian), RCD Publishers, 2005, 512 pp. www.rcd.ru/details/777

(4) J. Marche, On Kontsevich integral of torus knots, Université Paris 7 preprint, October 2003, arXiv: math.GT/0310111.

(5) J. Marche, A computation of Kontsevich integral of torus knots, Algebraic and Geometric Topology 4-51 (2004) 1155–1175, arXiv:math.GT/0404264.

(6) J. Marche, Surgery on a single clasper and the 2-loop part of the Kontsevich integral, Université Paris 7 preprint, October 2004, arXiv:math.GT/0410272.

(7) M. Mariño, Chern-Simons theory, matrix integrals, and perturbative three-manifold invariants, Harvard University HUTP-02/A029 preprint, arXiv:hep-th/0207096.

(8) J. F. Martins, Knot theory With the Lorentz group, University of Nottingham preprint, September 2003, arXiv:math.QA/0309162.

(9) J. F. Martins, On the analytic properties of the z-coloured Jones polynomial, University of Nottingham preprint, October 2003, arXiv:math.QA/0310394.

(10) G. Masbaum, Matrix-tree theorems and the Alexander-Conway polynomial, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 201-214, arXiv:math.CO/0211063.

(11) G. Masbaum and A. Vaintrob, Milnor numbers, spanning trees, and the Alexander-Conway polynomial , Université Paris 7 and University of Oregon preprint, November 2001, arXiv:math.GT/0111102.

(12) G. Massuyeau, Invariants de type fini des variétés de dimension trois et structures spinorielles, Ph.D. thesis, Université de Nantes, October 2002.

(13) G. Massuyeau, Cohomology rings, Rochlin function, linking pairing and the Goussarov-Habiro theory of 3-manifolds, Algebraic and Geometric Topology 3-41 (2003) 1139–1166, arXiv:math.GT/0307396.

(14) G. Massuyeau and J.-B. Meilhan, Characterization of Y_{2}-equivalence for homology cylinders, Jour. of Knot Theory and its Ramifications 12-4 (2003) 493–522, arXiv:math.GT/0203179.

(15) G. Massuyeau, Some finiteness properties for the Reidemeister-Turaev torsion of three-manifolds, arXiv:math.GT/0507214.

(16) J. Mattes, Functions on the moduli space of flat G_{2}-connections on a Riemann surface, to appear in Proc. Amer. Math. Soc.

(17) S. Matveev and M. Polyak, Cubic complexes and finite type invariants, Invariants of knots and 3manifolds (Kyoto 2001), Geometry and Topology Monographs 4 215-233, arXiv:math.GT/0204085.

(18) M. McDaniel, Decomposition of webs, web document, http://www.aquinas.edu/homepages/mcdanmic/decompos.htm, April 1999.

(19) M. McDaniel and Y. Rong, On the dimensions of Vassiliev invariants coming from link polynomials, George Washington University preprint, March 1997.

(20) J.-B. Meilhan, Invariants de type fini des cylindres d’homologie et des string links, Ph.D. thesis, Université de Nantes, December 2003.

(21) J.-B. Meilhan, Goussarov-Habiro theory for string links and the Milnor-Johnson correspondence, Université de Nantes preprint, February 2004, arXiv:math.GT/0402035.

(22) J.-B. Meilhan, On Vassiliev invariants of order two for string links, Jour. of Knot Theory and its Ramifications, v. 14, no. 5 (2005), 665–687. arXiv:math.GT/0402036.

(23) J.-B. Meilhan, Borromean surgery formula for the Casson invariant, arXiv:math.GT/0509335.

(24) J.-B. Meilhan, On surgery along Brunnian links in 3-manifolds, arXiv:math.GT/0603421.

(25) S. A. Melikhov, Colored finite type invariants and multi-variable analogue of the Conway polynomial, Steklov Mathematical Institute preprint, December 2003, arXiv:math.GT/0312007.

(26) S. A. Melikhov and D. Repovs, A geometric filtration of links modulo knots: II. Comparison, Moscow State University and University of Ljubliana preprint, March 2001, arXiv:math.GT/0103114.

(27) B. Mellor, The intersection graph conjecture for loop diagrams, Jour. of Knot Theory and its Ramifications 9(2) (2000) 187–211, arXiv:math.GT/9807033.

(28) B. Mellor, Finite type link homotopy invariants, Jour. of Knot Theory and its Ramifications 8(6) (1999) 773–787, arXiv:math.GT/9807162.

(29) B. Mellor, Finite type link homotopy invariants II: Milnor’s invariants, Jour. of Knot Theory and its Ramifications 9(6) (2000) 735–758, arXiv:math.GT/9812119.

(30) B. Mellor, Finite type link concordance invariants, Jour. of Knot Theory and its Ramifications 9(3) (2000) 367–385, arXiv:math.GT/9904169.

(31) B. Mellor, A few weight systems arising from intersection graphs, Michigan Math. Jour. 51-3 (2003) 509–536, arXiv:math.GT/0004080.

(32) B. Mellor, Intersection graphs for string links, Loyola Marymount University preprint, December 2003, arXiv:math.GT/0312347.

(33) B. Mellor, Tree diagrams for string links, Loyola Marymount University preprint, May 2004, arXiv: math.GT/0405537.

(34) B. Mellor, Weight Systems for Milnor Invariants, Loyola Marymount University preprint, January 2005, arXiv:math.GT/0501280.

(35) B. Mellor and D. Thurston, On the existence of finite type link homotopy invariants, Jour. of Knot Theory and its Ramifications 10(7) (2001) 1025–1040, arXiv:math.GT/0010206.

(36) P. M. Melvin and H. R. Morton, The coloured Jones function, Comm. Math. Phys. 169 (1995) 501–520.

(37) G. Meng, Bracket models for weight systems and the universal Vassiliev invariants, Hong Kong University for Science and Technology preprint, April 1994.

(38) A. B. Merkov, Finite-order invariants of ornaments, Merkov, A. B. (1998) Finite order invariants of ornaments, J. of Math. Sciences 90:4, 2215–2273.

(39) A. B. Merkov, Vassiliev invariants classify flat braids, in Differential and Symplectic Topology of Knots and Curves (S. Tabachnikov, ed.) AMS Translations Ser. 2 vol. 190 (1999), Providence.

(40) A. B. Merkov, Vassiliev invariants classify plane curves and doodles, Mat. Sbornik 194:9, 31-62.

(41) A.B.Merkov, On the Classification of Ornaments Advnces in Sov. Math., Vol. 21 (1994), 199-211.

(42) A.B.Merkov, Segment–arrow diagrams and invariants of ornaments Mat. Sbornik 191:11, 47-78 (2000).

(43) H. A. Miyazawa and A. Yasuhara, Classification of n-component Brunnian links up to C_{n}-move, Tsuda College and Tokyo Gakugei University preprint, December 2004, arXiv:math.GT/0412490.

(44) Iain Moffatt, The Aarhus integral and the mu-invariants, arXiv:math.GT/0511435.

(45) H. R. Morton, The coloured Jones function and Alexander polynomial for torus knots, Math. Proc. Camb. Philos. Soc. 117 (1995) 129–136.

(46) H. R. Morton and P. R. Cromwell, Distinguishing mutants by knot polynomials, J. Knot Theory Ramifications 5 (1996), no. 2, 225–238. (preprint version: July 1994).

(47) H. R. Morton and N. Ryder, Invariants of genus 2 mutants, arXiv:0708.0514.

(48) D. Moskovich, Framing and the self-linking integral, University of Tokyo preprint, November 2002, arXiv:math.QA/0211223. Far East J. of Math. Sci., vol. 14, no. 2, pp. 165–183 (2004).

(49) D. Moskovich, Acyclic Jacobi Diagrams. arXiv:math.GT/0507351.

(50) D. Moskovich, T. Ohtsuki. Vanishing of 3-Loop Jacobi Diagrams of Odd Degree. arXiv:math.GT/0511602.

(51) J. Mostovoy and T. Stanford, On a map from pure braids to knots, Universidad National Autónoma de México and United States Naval Academy preprint, July 1999. See also math.GT/9907088.

(52) J. Mostovoy and T. Stanford, On invariants of Morse knots, Universidad National Autónoma de México and United States Naval Academy preprint, August 2000, arXiv:math.GT/0008096.

(53) J. Mostovoy and S. Willerton, Free groups and finite type invariants of pure braids, Max-Planck-Institut preprint MPI-1999-54, May 1999. See also math.GT/9909070.

(54) H. Murakami, Vassiliev invariants of type two for a link, Proc. Amer. Math. Soc. 124 (1996), 3889–3896.

(55) H. Murakami, Calculations of the Casson-Walker-Lescop invariant from chord diagrams, in Knot theory (V. F. R. Jones, J. Kania-Bartoszyńska, J. H. Przytycki, P. Traczyk, and V. G. Turaev, eds.), Banach Center Publications 42 243–254, Warsaw 1998.

(56) H. Murakami, Hyperbolic three-manifolds with trivial finite type invariants, Waseda University and Mittag-Leffler Institute preprint, February 1999. See also math.GT/9902018.

(57) H. Murakami, A weight system derived from the multivariable Conway potential function, to appear in the J. London Math. Soc. See also math.GT/9903108.

(58) H. Murakami and J. Murakami, The colored Jones polynomials and the simplicial volume of a knot, Waseda University, Osaka University, and Mittag-Leffler Institute preprint, May 1999. See also math.GT/9905075.

(59) H. Murakami and T. Ohtsuki, Finite type invariants of knots via their Seifert matrices, Waseda University and Tokyo Institute of Technology preprint, December 1998. See also math.GT/9903069.

(60) J. Murakami, The Casson invariant for a knot in a 3-manifold, in Proc. of the Århus Conf. Geometry and physics, (J. E. Andersen, J. Dupont, H. Pedersen, and A. Swann, eds.), lecture notes in pure and applied mathematics 184 (1997) 459–469, Marcel Dekker, New-York.

(61) J. Murakami, Representations of the mapping class group via the universal perturbative invariant, in Proc. Knots 96 (S. Suzuki, ed.) (1997) 573–586, World Scientific. See correction in http://fuji.math.sci.osaka-u.ac.jp/˜ jun/papers/knots96c.html

(62) J. Murakami, On web diagrams, Osaka University preprint, January 1999.

(63) J. Murakami, Finite-type invariants detecting the mutant knots Proceedings of the Conference in Knot Theory dedicated to Professor Kunio Murasugi for his 70th birthday, Toronto July 13th17th, 1999, Toronto, March 2000, pp. 258–267.

(64) J. Murakami and T. Ohtsuki, Topological quantum field theory for the universal quantum invariant, to appear in Comm. Math. Phys.

(65) K. Murasugi, knot Theory and Its applications, Birkhäuser, Boston 1996.

4.14. References beginning with N.

(1) Y. Nakanishi and Y. Ohyama, Delta link homotopy for two component links, II, Jour. of Knot Theory and its Ramifications, to appear.

(2) G. Naot, On Chern-Simons theory with an inhomogeneous gauge group, University of Toronto preprint, October 2003, arXiv:math.GT/0310366.

(3) K. Y. Ng, Groups of ribbon knots, Topology 37 (1998) 441–458, arXiv:q-alg/9502017.

(4) K. Y. Ng, Addendum to “groups of ribbon knots”, arXiv:math.GT/0310074.

(5) K. Y. Ng and T. Stanford, On Gusarov’s groups of knots, Proc. Camb. Phil. Soc. 126-1 (1999) 63–76.

(6) D. Nguyen, A biaction of the braided category of representations of a quantum group and a Kontsevich integral, University of Ottawa preprint, January 1994.

(7) D. Nguyen, Equivariant Vassiliev invariants and a Kontsevich integral, University of Ottawa preprint, January 1994.

(8) H. Nishihara, On invariant symmetric 2-tensors of nilpotent Lie algebras of maximal rank, Osaka University preprint, October 2001.

(9) H. Nishihara, On weight systems derived from Heisenberg Lie algebras, Osaka University preprint, September 2002.

(10) H. Nishihara, The loop-degrees of Jacobi diagrams and Lie algebras, Osaka University preprint, March 2003.

(11) S. D. Noble and D. J. A. Welsh, A weighted graph polynomial from chromatic invariants of knots, Symposium à la Mémoire de François Jaeger (Grenoble, 1998), Ann. Inst. Fourier (Grenoble) 49-3 (1999) 1057–1087.

(12) T. Nowik, Finite order q-invariants of immersions of surfaces into 3-space, Math. Zeit. 236 (2001) 215–221, arXiv:math.GT/9903161.

(13) T. Nowik, Order one invariants of immersions, Math. Ann., to appear, arXiv:math.GT/0103157.

(14) T. Nowik, Higher order invariants of immersions of surfaces into 3-space, Bar Ilan University preprint, July 2002, arXiv:math.GT/0207087.

(15) T. Nowik, Formulae for order one invariants of immersions and embeddings of surfaces, Bar Ilan University preprint, May 2003, arXiv:math.GT/0305070.

(16) T. Nowik, Immersions of non-orientable surfaces, Bar Ilan University preprint, October 2003, arXiv: math.GT/0310025.

4.15. References beginning with O.

(1) T. Ochiai, The combinatorial Gauss diagram formula for Kontsevich integral, University of Tokyo preprint, June 2000, arXiv:math.GT/0006184.

(2) E. Ogasa, Ribbon-moves of 2-links preserve the μ-invariant of 2-links, University of Tokyo preprint, April 2000. See also math.GT/0004008.

(3) T. Ohtsuki, Finite type invariants of integral homology 3-spheres, Jour. of Knot Theory and its Ramifications 5(1) (1996) 101–115.

(4) T. Ohtsuki, On some invariants of 3-manifolds, in Proc. of the Århus Conf. Geometry and physics, (J. E. Andersen, J. Dupont, H. Pedersen, and A. Swann, eds.), lecture notes in pure and applied mathematics 184 (1997) 411–427, Marcel Dekker, New-York.

(5) T. Ohtsuki, Combinatorial quantum method in 3-dimensional topology, Math. Soc. of Japan Memoirs 3, Tokyo 1999.

(6) T. Ohtsuki, The perturbative SO(3) invariant of rational homology 3-spheres recovers from the universal perturbative invariant, to appear in Topology.

(7) T. Ohtsuki, A filtration of the set of integral homology 3-spheres, Tokyo Institute of Technology preprint, July 1999.

(8) T. Ohtsuki, Topological quantum field theory for the LMO invariant, Tokyo Institute of Technology preprint, July 1999.

(9) T. Ohtsuki, Quantum Invariants: A Study of Knots, 3-Manifolds, and their Sets, Series on Knots and Everything 29, World Scientific, Singapore 2002.

(10) T. Ohtsuki, A cabling formula for the 2-loop polynomial of knots, RIMS (Kyoto) preprint, October 2003, arXiv:math.GT/0310216.

(11) T. Ohtsuki, Problems on invariants of knots and 3-manifolds, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 377–572.

(12) Y. Ohyama, Vassiliev invariants and similarity of knots, Proc. Amer. Math. Soc. 123 (1995) 287–291.

(13) Y. Ohyama, Twisting of two strings and Vassiliev invariants, Topology Appl. 75 (1997) 201–215.

(14) Y. Ohyama, Remarks on C_{n}-moves for links and Vassiliev invariants of order n, Nagoya Institute of Technology preprint.

(15) Y. Ohyama, Web diagrams and realization of Vassiliev invariants by knots, Jour. of Knot Theory and Its Ramifications 9(5) (2000) 693–701.

(16) Y. Ohyama and K. Taniyama, Vassiliev invariants of knots in a spatial graph, Pacific Journal of Mathematics 200 (2001) 191–205.

(17) Y. Ohyama and T. Tsukamoto, On Habiro’s C_{n}-moves and Vassiliev invariants of order n, Jour. of Knot Theory and its Ramifications 8(1) (1999) 15–23.

(18) Y. Ohyama and H. Yamada, Delta and clasp-pass distance and Vassiliev invariants of knots, Nagoya Institute of Technology and Waseda University preprint, January 2001.

(19) M. Okamoto, Vassiliev invariants of type 4 for algebraically split links, Kobe Jour. of Math. 14-2 (1997) 145–196.

(20) M. Okamoto, On Vassiliev invariants for algebraically split links, Jour. of Knot Theory and its Ramifications 7-6 (1998) 807–835.

(21) P-Z. Ong and S-W. Yang, Cohomology of Vassiliev complex and passing product, National Taiwan University preprint, September 1994.

(22) P-Z. Ong and S-W. Yang, Homology of Vassiliev complex, National Taiwan University preprint, August 1995.

(23) P-Z. Ong and S-W. Yang, Higher degree Vassiliev invariants: a survey article, National Taiwan University preprint, September 1995.

(24) T. Ozawa, Finite order topological invariants of plane curves, Jour. of Knot Theory and its Ramifications 8(1) (1999) 33–47.

4.16. References beginning with P.

(1) P. Papi and C. Procesi (assisted by P. Cellini), Invarianti di nodi, survey monograph, to appear in Quaderni Dell’U.M.I.

(2) B. Patureau-Mirand, Caractères sur l’algèbre de diagrammes trivalents Λ, Geometry and Topology 6-20 (2002) 565–607, arXiv:math.GT/0107137.

(3) B. Patureau-Mirand, Invariant d’entrelacs associé à la représentation des spineurs de so(7), Université Paris 7 preprint, July 2001, arXiv:math.GT/0107138.

(4) B. Patureau-Mirand, Non injectivity of the “hair” map, Université Paris 7 preprint, February 2002, arXiv:math.GT/0202065.

(5) B. Patureau-Mirand, Quantum link invariant from the Lie superalgebra D(2, 1, α), Université de Bretagne-Sud preprint, April 2004, arXiv:math.GT/0404548.

(6) R. Picken, Knot invariants from a Kohno-Kontsevich integral for braids, q-alg/9707010 preprint, July 1997.

(7) S. Piunikhin, Weights of Feynman diagrams, link polynomials and Vassiliev knot invariants, Jour. of Knot Theory and its Ramifications 4(1) (1995) 165–188.

(8) S. Piunikhin, Combinatorial expression for universal Vassiliev link invariant, Comm. Math. Phys. 168-1 1–22.

(9) S. Piunikhin, Addendum to the paper “Combinatorial expression for universal Vassiliev link invariant, November 1993, preprint.

(10) S. Podkorytov, Vassiliev invariants of smooth submanifolds, Abstract of a talk (1998, in Russian).

(11) S. Podkorytov, On the maps of a sphere into a simply connected space, Zapiski POMI, v. 329 (2005), pp. 159–194 (in Russian).

(12) S. Poirier, Rationality results for the configuration space integral of knots, Université de Grenoble I preprint, December 1998. See also math.GT/9901028.

(13) S. Poirier, The limit configuration space integral for tangles and the Kontsevich integral, Université de Grenoble I preprint, February 1999, arXiv:math.GT/9902058.

(14) S. Poirier, The configuration space integral for links and tangles in ℝ^{3}, Algebraic and Geometric Topology, 2-40 (2002) 1001-1050, arXiv:math.GT/0005085.

(15) S. Poirier, A note on the configuration space integral, preprint, January 2002.

(16) M. Polyak, Link homotopy and Vassiliev knot invariants, IHES preprint, June 1996.

(17) M. Polyak, On Milnor’s triple linking number, Comp. Rend. A. S. Paris 325 Serie I (1997), 77–82.

(18) M. Polyak, Skein relations and Gauss diagram formulas for Milnor’s μ-invariants, Tel Aviv University preprint, December 1998.

(19) M. Polyak, On the algebra of arrow diagrams, Let. Math. Phys. 51 (2000) 275–291.

(20) M. Polyak, Feynman diagrams for pedestrians and mathematicians, Technion preprint, June 2004, arXiv:math.GT/0406251.

(21) M. Polyak and O. Viro, Gauss diagram formulas for Vassiliev invariants, Int. Math. Res. Notes 11 (1994) 445–454.

(22) M. Polyak and O. Viro, On the Casson knot invariant, Jour. of Knot Theory and its Ramifications, to appear, arXiv:math.GT/9903158.

(23) V. V. Prasolov and A. B. Sossinsky, Knots, links, braids and 3-manifolds: an introduction to the new invariants in low-dimensional topology, Translations of Mathematical Monographs 154, American Mathematical Society 1997.

(24) J. H. Przytycki, Vassiliev-Gusarov skein modules of 3-manifolds and criteria for periodicity of knots, in Low-Dimensional Topology, Knoxville 1992, 143–162, International Press, Cambridge MA, 1994.

(25) J. H. Przytycki, Symmetric knots and billiard knots, Chapter 20 of Ideal Knots, Series on Knots and Everything 19, Eds. A. Stasiak, V. Katrich and L. Kauffman, World Scientific (1999) 374–414, arXiv:math.GT/0405151.

(26) J. H. Przytycki, Graphs and links, arXiv:math.GT/0601227.

(27) J. H. Przytycki and A. S. Sikora, Topological insights from the Chinese rings, Proc. Amer. Math. Soc., to appear, arXiv:math.GT/0007134.

(28) J. H. Przytycki and K. Taniyama, The Kanenobu-Miyazawa conjecture and the Vassiliev-Gusarov skein modules based on mixed crossings, Proc. Amer. Math. Soc. 129 (2001) 2799–2802.

4.17. References beginning with Q.

(1) F. Quinn, Dual decompositions of 4-manifolds II: linear link invariants, Virginia Tech preprint, September 2001, arXiv:math.GT/0109148.

4.18. References beginning with R.

(1) R. Randell, Invariants of piecewise-linear knots, in Knot theory (V. F. R. Jones, J. Kania-Bartoszyńska, J. H. Przytycki, P. Traczyk, and V. G. Turaev, eds.), Banach Center Publications 42 307–319, Warsaw 1998.

(2) J. Roberts, Rozansky-Witten theory, University of California at San Diego preprint, arXiv:math.QA/0112209.

(3) J. Roberts and J. Sawon, Generalisations of Rozansky-Witten invariants, Invariants of knots and 3manifolds (Kyoto 2001), Geometry and Topology Monographs 4 263-279, arXiv:math.DG/0112210.

(4) J. Roberts, S. Willerton, On the Rozansky-Witten weight systems, arXiv:math.DG/0602653.

(5) P. Røgen, On density of the Vassiliev invariants, Jour. of Knot Theory and its Ramifications 8(2) (1999) 249–252.

(6) P. Røgen and B. Fain, Automatic classification of protein structure by using Gauss integrals, Proc. Natl. Acad. Sci. USA 100-1 (2003) 119–124.

(7) Y. Rong, Link polynomials of higher order, George Washington University preprint, January 1995.

(8) L. Rozansky, The trivial connection contribution to Witten’s invariant and finite type invariants of rational homology spheres, Comm. Math. Phys. 183 (1997) 23–54. See also q-alg/9503011.

(9) L. Rozansky, Witten’s invariants of rational homology spheres at prime values of K and the trivial connection contribution, Comm. Math. Phys. 180 (1996) 297–324. See also q-alg/9504015.

(10) L. Rozansky, On finite type invariants of links and rational homology spheres derived from the Jones polynomial and Witten-Reshetikhin-Turaev invariant, in Proc. of the Århus Conf. Geometry and physics, (J. E. Andersen, J. Dupont, H. Pedersen, and A. Swann, eds.), lecture notes in pure and applied mathematics 184 (1997) 379–397, Marcel Dekker, New-York. See also q-alg/9511025.

(11) L. Rozansky, Higher order terms in the Melvin-Morton expansion of the colored Jones polynomial, Comm. Math. Phys. 183 (1997) 291–306. See also q-alg/9601009.

(12) L. Rozansky, A rational structure of generating functions for Vassiliev invariants, Yale University preprint, July 1999.

(13) L. Rozansky, A rationality conjecture about Kontsevich integral of knots and its implications to the structure of the colored Jones polynomial, Proccedings of the Pacific Institute for the Mathematical Sciences Workshop (Calgary 1999), Topology and its Applications 127 (2003) 47–76, arXiv: math.GT/0106097.

(14) L. Rozansky, A universal U(1)-RCC invariant of links and rationality conjecture, University of North Carolina preprint, January 2002, arXiv:math.GT/0201139.

(15) L. Rozansky and E. Witten, Hyper-Kähler geometry and invariants of three-manifolds, Selecta Mathematica, New Series, 3 (1997) 401–458, arXiv:hep-th/9612216.

4.19. References beginning with S.

(1) J. Sawada, A fast algorithm for generating non-isomorphic chord diagrams, SIAM Jour. on Discrete Math., 15-4 (2002) 546–561.

(2) S. F. Sawin, Primitive Vassiliev invariants, MIT preprint, April 1993.

(3) S. F. Sawin, Finite-degree link invariants and connectivity, Jour. of Knot Theory and its Ramifications 5(1) (1996) 117–136.

(4) J. Sawollek, An orientation-sensitive Vassiliev invariant for virtual knots, Universität Dortmund preprint, March 200, arXiv:math.GT/0203123.

(5) J. Sawon, Rozansky-Witten invariants of hyperKähler manifolds, Ph.D. thesis, University of Cambridge, October 1999.

(6) J. Sawon, The Rozansky-Witten invariants of hyperKähler manifolds, Proceedings of the 7th International Conference on Differential Geometry and Applications, Brno, August 1998.

(7) J. Sawon, A new weight system on chord diagrams via hyperkähler geometry, Proceedings of the Second Meeting on Quaternionic Structures in Mathematics and Physics, Roma, September 1999, arXiv:math.DG/0002218.

(8) J. Sawon, Topological quantum field theory and hyperKḧler geometry, Proceedings of the 7th Gökova Geometry-Topology Conference and Turk. J. Math. 25 (2001) 169–194, arXiv:math.QA/0009222.

(9) J. Sawon, When is a Lie algebra not a Lie algebra? Proceedings of the IXth Oporto Meeting on Geometry, Topology & Physics, September-October 2000.

(10) R. Schneiderman, Algebraic linking numbers of knots in 3-manifolds, Algebraic and Geometric Topology 3-31 (2003) 921–968, arXiv:math.GT/0202024.

(11) R. Schneiderman, Simple Whitney towers, half-gropes and the Arf invariant of a knot, New York University preprint, October 2003, arXiv:math.GT/0310304.

(12) R. Schneiderman and P. Teichner, Whitney towers and the Kontsevich integral, Proceedings of the Casson fest, Geometry and Topology Monographs 7 101–134, arXiv:math.GT/0401441.

(13) K. Selwat, Niezmienniki Vassilieva krzywych na plaszczyznie, University of Wroclaw Master’s thesis, 1995.

(14) K. Selwat, The first order semilocal Vassiliev invariants of plane curves, Jour. of Knot Theory and its Ramifications 7(8) (1998).

(15) N. Shirokova, The space of 3-manifolds and Vassiliev finite-type invariants, University of Chicago preprint, May 1997. See also q-alg/9705008.

(16) N. Shirokova, On the classification of Floer-type theories, arXiv:0704.1330.

(17) A. Shumakovitch, Shadow formula for the Vassiliev invariant of degree two, Topology 36-2 (1997).

(18) D. J. Sinha, The topology of spaces of knots, Brown University preprint, February 2002, arXiv: math.AT/0202287.

(19) D. Sinha, Operads and Knot Spaces, preprint.

(20) S. A. Sirotine, Approximation of knot invariants by Vassiliev invariants, Rice University preprint, April 1995.

(21) C. T. Snydal, Vassiliev knot invariants arising from the Lie superalgebra gl(1 | 1), University of North Carolina physics undergraduate thesis, May 1995.

(22) E. Soboleva, Vassiliev knot invariants coming from Lie algebras and 4-invariants, Jour. of Knot Theory and its Ramifications 10(1) (2001).

(23) A. B. Sossinsky, Feynman diagrams and Vassiliev invariants, IHES preprint IHES/M/92/13, February 1992.

(24) T. Stanford, Finite type invariants of knots, links, and graphs, Topology 35-4 (1996) 1027–1050.

(25) T. Stanford, Braid commutators and Vassiliev invariants, Pacific Jour. of Math. 174-1 (1996).

(26) T. Stanford, The functoriality of Vassiliev-type invariants of links, braids, and knotted graphs, Jour. of Knot Theory and its Ramifications 3(3) (1994) 247–262.

(27) T. Stanford, Computing Vassiliev’s invariants, Topology and its Applications 77-3 (1997). See also programs at ftp://geom.umn.edu/pub/contrib/vassiliev.tar.Z.

(28) T. Stanford, Vassiliev invariants and knots modulo pure braid subgroups, United States Naval Academy preprint, May 1998, arXiv:math.GT/9805092.

(29) T. Stanford, Four observations on n-triviality and Brunnian links, United States Naval Academy preprint, July 1998. See also math.GT/9807161.

(30) T. Stanford, Braid commutators and delta finite-type invariants, United States Naval Academy preprint, July 1999. See also math.GT/9907071.

(31) T. Stanford, Some computational results on mod 2 finite-type invariants of knots and string links, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 363–376.

(32) T. Stanford and R. Trapp, On knot invariants which are not of finite type, United States Naval Academy and California State University (San Bernardino) preprint, March 1999. See also math.GT/9903057.

(33) A. Stoimenow, On Harrison cohomology and a conjecture by Drinfel’d, diploma thesis, Humboldt University, Berlin, 1995 (also available in German as Über Harrison-kohomologie und die Drinfel’dvermutung).

(34) A. Stoimenow, Stirling numbers, Eulerian idempotents and a diagram complex, version 27/07/2005, published in J. Of Knot Theory and Its Ram. 7(2) (1998), 231-256.

(35) A. Stoimenow, On the number of chord diagrams, Discr. Math. 218 (2000), 209–233. (based on a preprint, January 1996).

(36) A. Stoimenow, Enumeration of chord diagrams and an upper bound for Vassiliev invariants, Jour. of Knot Theory and its Ramifications 7(1) (1998) 94–114.

(37) A. Stoimenow, On enumeration of chord diagrams and asymptotics of Vassiliev invariants, PhD thesis, Humboldt University, Berlin, 1999.

(38) A. Stoimenow, Positive knots, closed braids and the Jones polynomial, Ann. Scuola Norm. Sup. Pisa Cl. Sci. 2(2) (2003), 237–285. (previous versions: Humboldt University preprint, May 1997, and math.GT/9805078).

(39) A. Stoimenow, Gauss sum invariants, Vassiliev invariants and braiding sequences, J. Of Knot Theory and Its Ram. 9(2) (2000), 221–269 (first appeared as a preprint, Humboldt University, May 1997).

(40) A. Stoimenow, The braid index and the growth of Vassiliev invariants Jour. of Knot Theory and its Ramifications, 8(6) (1999), 799–813.

(41) A. Stoimenow, Vassiliev invariants on fibered and mutually obverse knots, Jour. of Knot Theory and its Ramifications 8(4) (1999) 511–519.

(42) A. Stoimenow, On finiteness of Vassiliev invariants and a proof of the Lin-Wang conjecture via braiding polynomials, J. Of Knot Theory and Its Ram. 10(5) (2001), special volume for the proceedings of the International Conference on Knot Theory ‘Knots in Hellas, 98”, 769–780.

(43) A. Stoimenow, Genera of knots and Vassiliev invariants, Jour. of Knot Theory and its Ramifications 8(2) (1999) 253–259.

(44) A. Stoimenow, Gauss sums on almost positive knots, math.GT/9803073. Compositio Mathematica 140(1) (2004), 228–254.

(45) A. Stoimenow, Mutant links distinguished by degree 3 Gauss sums, (online version: 27/04/1998), Proceedings of the International Conference on Knot Theory ”Knots in Hellas, 98”, Series on Knots and Everything 24, World Scientific, 2000.

(46) A. Stoimenow, Some minimal degree Vassiliev invariants not realizable by the HOMFLY and Kauffman polynomial, math.GT/9807076. C. R. Acad. Bulgare Sci. 54(4) (2001), 9–14.

(47) A. Stoimenow, Knots of genus one, Proc. Amer. Math. Soc. 129(7) (2001), 2141–2156 (first appeared as a preprint, Munich, 1998).

(48) A. Stoimenow, A survey on Vassiliev invariants for knots, Mathematics and Education in Mathematics, Proceedings of 27th Spring Conference of the Union of Bulgarian Mathematicians, Pleven 1998.

(49) A. Stoimenow, Vassiliev invariants and rational knots of unknotting number one, math.GT/9909050. Topology 42(1) (2003), 227–241.

(50) A. Stoimenow, Polynomial and polynomially growing knot invariants, Preprint, current version 28/06/2005.

(51) A. Stoimenow, The Conway Vassiliev invariants on twist knots, Kobe J. Math. 16(2) (1999), 189-193.

(52) A. Stoimenow, On cabled knots and Vassiliev invariants (not) contained in knot polynomials, to appear in Canadian J. Math.

(53) A. Stoimenow, Some applications of Tristram-Levine signatures, Advances in Mathematics 194(2) (2005), 463-484.

(54) Y. Suetsugu, Kontsevich invariant for links in a donut and links of satellite form, Osaka J. of Math. 33 (1996) 823–828.

4.20. References beginning with T.

(1) S. Tabachnikov. Invariants of triple points free plane curves. J. of Knot Theory and its Ramif., 5 (1996), 531-552.

(2) T. Takamuki, The Kontsevich invariant and relations of multiple zeta values, Kobe J. Math. 16 (1999), 27-43.

(3) K. Taniyama and A. Yasuhara, Band description of knots and Vassiliev invariants, to appear in Mathematical Proceedings of the Cambridge Philosophical Society, arXivmath.GT/0003021.

(4) K. Taniyama and A. Yasuhara, Local moves on spatial graphs and finite type invariants, Pac. Jour. of Math., to appear, arXiv:math.GT/0106173.

(5) P. Teichner, Knots, von Neumann signatures, and grope cobordism, Proceedings of the ICM, Beijing 2002, 2 437–446, arXiv:math.GT/0304299.

(6) G. Thompson, Holomorphic vector bundles, knots and the Rozansky-Witten invariants, ICTP Trieste preprint IC/2000/9, February 2000. See also hep-th/0002168.

(7) D. Thurston, Integral expressions for the Vassiliev knot invariants, Harvard University senior thesis, April 1995, arXiv:math.QA/9901110.

(8) D. Thurston, Wheeling: A diagrammatic analogue of the Duflo isomorphism, University of California at Berkeley Ph.D. thesis, May 2000, arXiv:math.QA/0006083.

(9) D. Thurston, The algebra of knotted trivalent graphs and Turaev’s shadow world, Invariants of knots and 3-manifolds (Kyoto 2001), Geometry and Topology Monographs 4 337–362, arXiv: math.GT/0311458.

(10) R. Trapp, Twist sequences and Vassiliev invariants, Jour. of Knot Theory and its Ramifications 3(3) (1994) 391–405.

(11) V. Turchin, Homology isomorphism of the complex of 2-connected graphs and the graph-complex of trees, in Topics in quantum groups and finite-type invariants, (B. Feigin and V. Vassiliev, eds.) AMS Translations Ser. 2 vol. 185 (1998) 147–153, Providence.

(12) V. Tourtchine, Sur l’homologie des espaces de nœuds non-compacts, Independent Moscow University and Université de Paris 7 preprint, October 2000, arXiv:math.QA/0010017.

(13) V. Tourtchine, On the homology of the spaces of long knots, Independent Moscow University and Université de Paris 7 preprint, May 2001, arXiv:math.QA/0105140.

(14) V. Tourtchine, Sur les questions combinatoires de la théorie spectrale des nœuds, PhD thesis, Université Paris 7, May 2002.

(15) V. Tourtchine, On the other side of the bialgebra of chord diagrams, Independent Moscow University and Université Catholique de Louvain preprint, November 2004, arXiv:math.QA/0411436.

(16) V. Turchin, Calculus of the first non-trivial 1-cocycle of the space of long knots, Independent University of Moscow preprint, February 2005, arXiv:math.AT/0502518.

(17) V. Tourtchine, What is 1-term relation for higher homology of long knots, arXiv:math.AT/0510667.

(18) V. Tourtchine, Hodge decomposition in the homology of long knots, arXiv:0812.0204.

(19) S. D. Tyurina, On formulas of the Lannes and Viro-Polyak type for the finite degree invariants, Mat. Zametki 4(66) (1999), arXiv:math.AT/9905160.

(20) S. D. Tyurina, Explicit formulas for the Vassiliev knot invariants, Kolomensky Pedagogical Institute preprint, May 1999. See also math.AT/9905161.

(21) S. D. Tyurina, Diagram formulas of the Viro-Polyak type for finit degree invariants, Russian Mathematical Surveys 54(3) (1999) 658–659, arXiv:math.AT/9905162.

(22) S. D. Tyurina, The Kontsevich integral for (2, n)-type torus knots, Proc. Rokhlin Memorial conference, Euler Institute, Saint-Petersburg, August 1999.

(23) S. D. Tyurina and A. Varchenko, A remark on sl_{2} approximation of the Kontsevich integral of the unknot, Zapiski Nauchnyh Seminarov POMI 299 30–37, arXiv:math.AT/0111201

(24) S. D. Tyurina and A. Varchenko, Finite order invariants for (n, 2)-torus knots and the curve Y^{2} = X^{3} + X^{2}, University of North Carolina at Chapel Hill preprint, February 2004, arXiv: math.GT/0402185.

4.21. References beginning with U.

4.22. References beginning with V.

(1) A. Vaintrob, Vassiliev knot invariants and Lie S-algebras, Math. Res. Lett. 1 (1994), no. 5, 579–595.

(2) A. Vaintrob, Universal weight systems and the Melvin-Morton expansion of the colored Jones knot invariant, q-alg/9605003 and University of Utah preprint, May 1996. (published in Russian in Itogi Nauki i Tehniki, t. 34, VINITI, 2001, pp. 195–216).

(3) A. Vaintrob, Melvin-Morton conjecture and primitive Feynman diagrams, q-alg/9605028 and University of Utah preprint, May 1996.

(4) V. A. Vassiliev, Cohomology of knot spaces, in Theory of Singularities and its Applications (Providence) (V. I. Arnold, ed.), Amer. Math. Soc., Providence, 1990.

(5) V. A. Vassiliev, Complements of discriminants of smooth maps: topology and applications, Trans. of Math. Mono. 98, Amer. Math. Soc., Providence, 1992; 2nd edition: 1994.

(6) V. A. Vassiliev, Invariants of knots and complements of discriminants, Developments in Mathematics, the Moscow School (V. I. Arnold and M. Monastyrsky, eds.), Chapman & Hall, London 1993.

(7) V. A. Vassiliev, On the spaces of polynomial knots, Matem. Sbornik, 187:2 (1996), 37–58.

(8) V. A. Vassiliev, On invariants and homology of spaces of knots in arbitrary manifolds, in Topics in quantum groups and finite-type invariants, (B. Feigin and V. Vassiliev, eds.) AMS Translations Ser. 2 vol. 185 (1998) 155–182, Providence.

(9) V. A. Vassiliev, Homology of i-connected graphs and invariants of knots, plane arrangements, etc., in The Arnoldfest, proceedings of a conference in honour of V. I. Arnold for his sixtieth birthday (E. Bierstone, B. Khesin, A. Khovanskii and J. E. Marsdeneds.), Fields Inst. Comm. 24 (1999) 451–469.

(10) V. A. Vassiliev, On combinatorial formulas for cohomology of spaces of knots, Moscow Math. J., 2001, 1:1, 91–123.

(11) V. A. Vassiliev, First degree invariants for spatial self-intersecting curves, Izvestiya Russian Ac. Sci., 2005, to appear.

(12) V. A. Vassiliev, Topological order complexes and resolutions of discriminant sets, Publications de l’Institut Mathématique Belgrade, Nouvelle série, t. 66(80), 1999, 165–185.

(13) V. A. Vassiliev, Homology of spaces of knots in any dimensions, Philos. Trans. London Royal Society A, 359:1784 (2001), 1343–1364.

(14) V. A. Vassiliev, Combinatorial computation of combinatorial formulas for knot invariants, Transact. of Moscow Math. Society 66 (2005), 3–94.

(15) V. A. Vassiliev, Knot invariants and singularity theory, In: Singularity Theory (Trieste, 1991), Le D.T., K. Saito and B. Teissier, eds., World Sci. Publishing, River Edge, NJ, 1995, 904–919. (Proceedings of the Colloquium in Singularity Theory, Trieste, ICTP, Aug. 19–Sep. 06, 1991).

(16) V. A. Vassiliev, Invariants of ornaments, Singularities and Bifurcations (V.I. Arnold, ed.), Advances in Soviet Math., vol. 21 (1994), 225–262 (AMS, Providence, R.I.). Online version.

(17) V. A. Vassiliev, Topology of two-connected graphs and homology of spaces of knots, In: ”Differential and symplectic topology of knots and curves,” AMS Transl. Ser. 2. Vol. 190. (S.L. Tabachnikov, ed.). AMS, Providence RI, 1999, 253–286.

(18) V. A. Vassiliev, On finite-order invariants of triple points free plane curves, In: AMS Transl. Ser. 2. Vol. 194. Volume dedicated to the 60-th birthday of D.B. Fuchs, (A. Astashkevich and S. Tabachnikov, eds.) AMS, Providence RI, 1999, 275–300. Online version.

(19) V. A. Vassiliev, Algorithms for the combinatorial realization of cohomology classes of spaces of knots, in: Fundamental Mathematics Today. In honor of the 10th anniversary of the Independent University of Moscow (S.Lando and O.Sheinman, eds.). Moscow, MCCME, 2003, pp. 10–31.

(20) V. A. Vassiliev, Combinatorial formulas for cohomology of spaces of knots, in: Advances in Topological Field Theory (J.Bryden, ed.) Springer-Verlag, 2004.

(21) V. A. Vassiliev, Complexes of connected graphs, The I.M. Gel’fand’s mathematical seminars 1990– 1992, L. Corvin, I. Gel’fand, J. Lepovsky, Eds.; Birkhäuser, Basel, 1993, p. 223–235.

(22) V. A. Vassiliev, Resolutions of discriminants and topology of their complements, in: New Developments in Singularity Theory, D. Siersma, C.T.C. Wall and V.M. Zakalyukin, eds., Kluwer Academic Publ., Dorderecht, 2001, 87–115.

(23) V. A. Vassiliev, Holonomic links and Smale principles for multisingularities. Jour. of Knot Theory and its Ramifications 6 (1) (1997) 115–123.

(24) V. V. Vershinin, On Vassiliev invariants for links in handlebodies, Jour. of Knot Theory and its Ramifications 7 (5) (1998) 701–712. See also q-alg/9709017.

(25) O. Viro, Self-linking number of a real algebraic link, Uppsala University preprint, July 1995, arXiv:alggeom/9410030.

(26) O. Viro, Encomplexing the writhe, Uppsala University and POMI St. Petersburg preprint, May 2000, arXiv:math.AG/0005162.

(27) P. Vogel, Invariants de Vassiliev des nœuds [d’après D. Bar-Natan, M. Kontsevich et V. A.Vassiliev], Séminaire Bourbaki 761 (1993) 1–17 (also, Asterisque 216 (1993) 213–232).

(28) P. Vogel, Algebraic structures on modules of diagrams, Université Paris VII preprint, July 1995 (revised 1997). Published (finally!) in J. Pure Appl. Algebra 215 (2011), no. 6, 1292–1339.

(29) P. Vogel, Vassiliev theory, Université Paris VII preprint, July 1999.

(30) P. Vogel, The universal Lie algebra, Université Paris VII preprint, June 1999.

(31) P. Vogel, Invariants de type fini, in Nouveaux invariants en géométrie et en topologie (François Dumas, Jean-Yves Le Dimet and Sylvie Paycha, eds.), Panor. Synthèses 11 99–128, Soc. Math. France, Paris 2001.

(32) P. Vogel, Vassiliev theory and the universal Lie algebra, Université Paris VII preprint, 2000.

(33) I. Volic, Finite type knot invariants and calculus of functors, Ph.D. thesis, Brown University, 2003, arXiv:math.AT/0401440.

(34) I. Volic, Configuration space integrals and Taylor towers for spaces of knots, University of Virginia preprint, January 2004, arXiv:math.GT/0401282.

(35) I. Volic, A survey of Bott-Taubes integration, University of Virginia preprint, February 2005, arXiv: math.GT/0502295.

(36) I. Volic, Calculus of the embedding functor and spaces of knots, arXiv:math.AT/0601268.

4.23. References beginning with W.

(1) S. Willerton, Vassiliev invariants and the Hopf algebra of chord diagrams, Math. Proc. Camb. Phil.

Soc. 119 (1996) 55–65.

(2) S. Willerton, A topological tie-in, Nature 368 No. 6467 (1994) 103–104.

(3) S. Willerton, A combinatorial half-integration from weight system to Vassiliev knot invariant, Jour.

of Knot Theory and its Ramifications 7(4) (1998) 519–526.

(4) S. Willerton, Vassiliev invariants as polynomials, in Knot theory (V. F. R. Jones, J. Kania-Bartoszyńska, J. H. Przytycki, P. Traczyk, and V. G. Turaev, eds.), Banach Center Publications 42 457–463, Warsaw 1998.

(5) S. Willerton, On the Vassiliev invariants for knots and for pure braids, Ph.D. thesis, University of Edinburgh.

(6) S. Willerton, Fishing with Vassiliev invariants, web document, http://www.sheffield.ac.uk/simonwillerton/fishing.html.

(7) S. Willerton, The Kontsevich integral and algebraic structures on the space of diagrams, Knots in Hellas ’98, Knots and Everything 24, World Scientific 2000, 530–546, arXiv:math.GT/9909151.

(8) S. Willerton, On the first two Vassiliev invariants, IRMA Strasbourg preprint, March 1999, arXiv:

math.GT/0104061. Experimental Mathematics, Vol. 11 (2002), No. 2, p. 289–296.

(9) S. Willerton, An almost-integral universal Vassiliev invariant of knots, Algebraic and Geometric Topology 2-29 (2002) 649–664, arXiv:math.GT/0105190.

4.24. References beginning with X.

4.25. References beginning with Y.

(1) H. Yamada, Delta distance and Vassiliev invariants of knots, Jour. of Knot Theory and its Ramifications 9(7) (2000) 967–974.

(2) M. Yamamoto, First order semi-local invariants of stable maps of 3-manifolds into the plane, Proc.

London Math. Soc. (3) 92 (2006), no. 2, 471–504.

(3) S.-W. Yang, Feynman integral, knot invariant and degree theory of maps, National Taiwan University preprint, September 1997, arXiv:q-alg/9709029.

(4) A. Yasuhara, C_{k}-moves on spatial theta-curves and Vassiliev invariants, George Washington University preprint, April 2001, arXiv:math.GT/0104177.

(5) A. Yasuhara, Brunnian local moves of knots and Vassiliev invariants, Tokyo Gakugei University preprint, December 2004, arXiv:math.GT/0412489.

(6) D. N. Yetter, Braided deformations of monoidal categories and Vassiliev invariants Kansas State University preprint, October 1997. See also q-alg/9710010.

(7) D. N. Yetter, Invariants of smooth 4-manifolds via Vassiliev theory. arXiv:math.GT/0510448.

4.26. References beginning with Z.

(1) D. Zagier, Vassiliev invariants and a strange identity related to the Dedekind eta-function, Topology 40(5) (2001) 945–960.

(2) V. Zapolsky, A functional characterization of Vassiliev knot invariants, PDMI preprint (2006, in Russian).

(3) J. Zhu, On singular braids, Jour. of Knot Theory and its Ramifications 6(3) (1997) 427–440.

(4) J. Zhu, On Jones knot invariants and Vassiliev invariants, New Zealand Jour. of Math., to appear.